### v2

# Towards Trustworthy Deep Neural Network through Euclidean Distance

Abstract
Uncertainty is a fundamental aspect of real-world scenarios, where perfect information is rarely available. Humans naturally develop complex internal models to navigate incomplete data and effectively respond to unforeseen or partially observed events. In machine learning, Focal Loss is commonly used to reduce misclassification rates by emphasizing hard-to-classify samples. However, it does not guarantee well-calibrated predicted probabilities and may result in models that are overconfident or underconfident. High calibration error indicates a misalignment between predicted probabilities and actual outcomes, affecting model reliability. This research introduces a novel loss function called Focal Calibration Loss (FCL), designed to improve probability calibration while retaining the advantages of Focal Loss in handling difficult samples. By minimizing the Euclidean norm through a strictly proper loss, FCL penalizes the instance-wise calibration error and constrains bounds. We provide theoretical validation for proposed method and apply it to calibrate CheXNet for potential deployment in web-based healthcare systems. Extensive evaluations on various models and datasets demonstrate that our method achieves SOTA performance in both calibration and accuracy metrics.
Machine Learning, ICML
1 Introduction
In critical applications such as medical imaging and diagnostic systems, well-calibrated models are paramount for ensuring accurate and trustworthy outcomes. By aligning predicted probabilities with the true likelihood of events, well-calibrated predictors reduce the discrepancy between a model’s confidence and actual results, fostering user trust and minimizing the need for post-processing calibration methods (e.g., temperature scaling (Guo et al., 2017) or isotonic regression (Niculescu-Mizil & Caruana, 2005)). However, while Focal Loss (Lin et al., 2017) effectively addresses class imbalance, it often leads to overly confident predictions (Charoenphakdee et al., 2021). This shortfall highlights the post-processing gap—the difference between raw model outputs and the properly calibrated probabilities required for reliable decision-making.
Figure 1: Grad-CAM (Selvaraju et al., 2017) Heatmap Visualization of CheXNet Model (Rajpurkar et al., 2017) Predictions for Pneumonia Detection. This figure shows Grad-CAM heatmaps from the CheXNet model applied to two ChestX-ray14 test cases for pneumonia detection. The model predicts the probability of each thoracic disease and generates a corresponding pathology likelihood map. Subfigures (1.a, 2.a) show outputs from the model trained with cross-entropy loss; (1.b, 2.b) with focal loss; and (1.c, 2.c) with focal calibration loss, illustrating the impact of different loss functions on localizing disease-relevant regions.
Motivation and Contributions. Recent studies have further established focal loss’s theoretical foundation in classification calibration (Charoenphakdee et al., 2021), showing that a classifier trained with focal loss can achieve the optimal Bayes classifier in theory. Nevertheless, focal loss is not a proper scoring rule, and training with it can yield underconfident or overconfident estimators—unreliable for class-posterior probability estimation. Proper scoring rules, such as the Brier score (Brier, 1950) or log-loss, encourage models to produce unbiased class-posterior probabilities (Lakshminarayanan et al., 2017) and ensure that post-processing does not significantly reduce the loss (Blasiok et al., 2024). Building on these insights, our work investigates whether integrating a proper loss into focal loss can improve its calibration as a class-posterior probability estimator. Our key contributions are:
  • Focal Calibration Loss (FCL): We propose a new loss function that combines the traditional Focal Loss with a proper calibration term, uniquely aligning predicted probabilities with true class-posterior probabilities to mitigate overconfidence and underconfidence.
  • Theoretical Guarantees: We prove that minimizing FCL yields superior probability and classification calibration compared to Focal Loss alone, effectively reducing the post-processing gap.
  • Empirical Validation: Through extensive experiments on both in-distribution and out-of-distribution image and NLP datasets, we demonstrate that FCL significantly enhances model calibration and classification performance by using the smooth calibration error further ensures calibration assessment. We also show that FCL improves anomaly localization, recognition, and robustness in Chest X-ray imaging applications.
2 Preliminaries
2.1 Problem Formulation
In multiclass classification, let $\mathcal{X}$ be the input space and $\mathcal{Y}=\{1,2,\ldots,K\}$ the label space of $K$ classes. [^1] Vectors are in bold (e.g., $\boldsymbol{x}$), scalars are in regular font (e.g., $x$), and the $i$-th element of a vector $\boldsymbol{x}$ is $x_{i}$. The indicator function is $\mathbf{1}_{[\cdot]}$, and $\boldsymbol{x}^{\top}$ denotes the transpose of $\boldsymbol{x}$. We observe a dataset $\mathcal{D}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$ drawn i.i.d. from an unknown distribution with density $p(\boldsymbol{x},y)$ on $\mathcal{X}\times\mathcal{Y}$. We seek a classifier $f:\mathcal{X}\to\mathcal{Y}$ that minimizes the classification risk (Charoenphakdee et al., 2021):
$\mathcal{R}^{\ell}(f)=\mathbb{E}_{(\boldsymbol{x},y)\sim p(\boldsymbol{x},y)}\bigl{[}\ell\bigl{(}f(\boldsymbol{x}),y\bigr{)}\bigr{]},$ (1)
where $\ell(\cdot,\cdot)$ is the zero-one loss, $\ell(f(\boldsymbol{x}),y)=\mathbf{1}_{[f(\boldsymbol{x})\neq y]}$. Although the true class-posterior probability vector $\boldsymbol{\eta}(\boldsymbol{x})=\bigl{[}\eta_{1}(\boldsymbol{x}),\ldots,\eta_{K}(\boldsymbol{x})\bigr{]}^{\top},\text{where }\eta_{k}(\boldsymbol{x})=\mathbb{P}\bigl{(}y=k\,\mid\,\boldsymbol{x}\bigr{)},$ would allow direct identification of the Bayes-optimal classifier:
$f^{\ell,*}(\boldsymbol{x})=\arg\max_{y\in\mathcal{Y}}\eta_{y}(\boldsymbol{x}),$ (2)
in practice we do not observe $\boldsymbol{\eta}(\boldsymbol{x})$ but only empirical samples $(\boldsymbol{x}_{i},y_{i})$. Hence, $\boldsymbol{\eta}(\boldsymbol{x})$ is purely theoretical and not assumed to be known.
2.2 Empirical Surrogate Risk
Minimizing $\mathcal{R}^{\ell}(f)$ in Eq. (1) directly is intractable because the zero-one loss is non-convex and non-differentiable (Bartlett et al., 2006; Zhang, 2004). Instead, one typically adopts a differentiable surrogate loss. Let
$f(\boldsymbol{x})=\hat{\boldsymbol{p}}(\boldsymbol{x})=\bigl{[}\hat{p}_{1}(\boldsymbol{x}),\ldots,\hat{p}_{K}(\boldsymbol{x})\bigr{]}^{\top}\in\Delta^{K},$
where $\Delta^{K}:=\{\boldsymbol{p}\in[0,1]^{K}\mid\sum_{k=1}^{K}p_{k}=1\}$. A common practice is to train a neural network with a softmax layer to output these probabilities, then define its prediction as:
$f^{*}(\boldsymbol{x})=\arg\max_{y\in\mathcal{Y}}\hat{p}_{y}(\boldsymbol{x}).$ (3)
The empirical surrogate risk is  (Vapnik et al., 1998):
$\hat{\mathcal{R}}^{\ell}(f)=\frac{1}{n}\sum_{i=1}^{n}\ell\bigl{(}f(\boldsymbol{x}_{i}),\mathbf{e}_{y_{i}}\bigr{)},$ (4)
where $\mathbf{e}_{y_{i}}\in\{0,1\}^{K}$ is the one-hot encoding of $y_{i}$. By choosing a suitable surrogate $\ell$, we obtain a differentiable training objective that, in theory, aligns with minimizing the zero-one loss (Bartlett et al., 2006; Vapnik et al., 1998).
2.3 Overconfidence (OC) and Underconfidence (UC)
Definition 2.1 (Overconfidence and Underconfidence).
Let $f(\boldsymbol{x})=\hat{\boldsymbol{p}}(\boldsymbol{x})$ be a model’s predicted probability vector, and let $\boldsymbol{\eta}(\boldsymbol{x})$ be the theoretical posterior. The model $f$ is said to be OC/UC at $\boldsymbol{x}\in\mathcal{X}$ if the highest predicted probability exceeds/less than the highest true probability:
$\delta_{\mathrm{OC}}(\boldsymbol{x})$ $:=\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\;>\;0,$ (5)
$\delta_{\mathrm{UC}}(\boldsymbol{x})$ $:=\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\;<\;0,$ (6)
2.4 Calibration Error
Calibration assesses whether the predicted probabilities align with empirical frequencies. Traditional binning-based metrics, such as Expected Calibration Error (ECE), can be sensitive to bin sizes. Therefore, an alternative is the Smooth Calibration Error (smCE) (Blasiok et al., 2024):
Definition 2.2 (Smooth Calibration Error (smCE) [^2] In real-world settings, the true posterior $\boldsymbol{\eta}(\boldsymbol{x})$ is unknown; hence calibration metrics must be estimated from empirical data $\{(\boldsymbol{x}_{i},y_{i})\}$ rather than from $\boldsymbol{\eta}(\boldsymbol{x})$.).
Let $\mathcal{H}$ be the class of all 1-Lipschitz functions $\phi:[0,1]\to[-1,1]$. For $\hat{\boldsymbol{p}}(\boldsymbol{x})=(\hat{p}_{1}(\boldsymbol{x}),\ldots,\hat{p}_{K}(\boldsymbol{x}))$ and one-hot $\boldsymbol{y}=(y_{1},\ldots,y_{K})$, the smooth Calibration Error is defined as:
$\text{smCE}(f):=\sup_{\phi\in\mathcal{H}}\;\mathbb{E}_{(\boldsymbol{x},y)\sim\mathcal{D}}\Bigl{[}\sum_{k=1}^{K}\bigl{(}y_{k}-\hat{p}_{k}(\boldsymbol{x})\bigr{)}\,\phi\bigl{(}\hat{p}_{k}(\boldsymbol{x})\bigr{)}\Bigr{]}.$ (7)
2.5 Focal Loss Bound
A popular surrogate is the focal loss (Lin et al., 2017), which addresses class imbalance by downweighting easy examples. In prior theoretical analyses (Mukhoti et al., 2020a; Charoenphakdee et al., 2021), focal loss can be related to the KL divergence from $\boldsymbol{\eta}(\boldsymbol{x})$ to $\hat{\boldsymbol{p}}(\boldsymbol{x})$, giving a bound of form:
$\mathcal{L}_{\text{focal}}\bigl{(}\hat{\boldsymbol{p}}(\boldsymbol{x}),\mathbf{e}_{y}\bigr{)}\geq\mathrm{KL}\bigl{(}\boldsymbol{\eta}(\boldsymbol{x})\,\big{\|}\,\hat{\boldsymbol{p}}(\boldsymbol{x})\bigr{)}+\mathbb{H}\bigl{[}\boldsymbol{\eta}(\boldsymbol{x})\bigr{]}-\gamma\,\mathbb{H}\bigl{[}\hat{\boldsymbol{p}}(\boldsymbol{x})\bigr{]},$ (8)
where $\mathbb{H}[\cdot]$ is the entropy function and $\gamma$ the focal “focusing” parameter. In this bound, $\boldsymbol{\eta}(\boldsymbol{x})$ appears only in a theoretical sense to illustrate how focal loss penalizes higher-entropy predictions more strongly as $\gamma$ increases. Importantly, no step in the focal-loss training algorithm relies on direct knowledge of $\boldsymbol{\eta}(\boldsymbol{x})$; the model is trained on empirical data using a focal-loss objective, while $\boldsymbol{\eta}(\boldsymbol{x})$ remains a theoretical concept.
2.6 Post-Processing Gap
Finally, the post-processing gap quantifies how much a smooth recalibration function $\kappa(\cdot)$ could improve a classifier’s calibration under a given loss. Formally:
Definition 2.3 (Post-Processing Gap [^3] A small post-processing gap suggests that the raw outputs of $f(\boldsymbol{x})$ are already near-optimal for calibration under smooth transformations. By contrast, a large gap indicates substantial room for improvement via recalibration techniques. (Blasiok et al., 2024)).
Let $\mathcal{K}$ be 1-Lipschitz mappings $\kappa:[0,1]\to[0,1]$. Define $\delta(v)=\kappa(v)-v$. For a chosen calibration loss $\mathcal{L}$ and predictor $f$, $\text{pGap}_{\mathcal{D}}(f)$ is defined as:
$\mathbb{E}_{(\boldsymbol{x},y)\sim\mathcal{D}}\bigl{[}\mathcal{L}\bigl{(}f(\boldsymbol{x}),y\bigr{)}\bigr{]}\;-\;\inf_{\kappa\in\mathcal{K}}\mathbb{E}_{(\boldsymbol{x},y)\sim\mathcal{D}}\bigl{[}\mathcal{L}\bigl{(}\kappa(f(\boldsymbol{x})),y\bigr{)}\bigr{]}.$ (9)
Reconciling Theory and Practice.
Although $\boldsymbol{\eta}(\boldsymbol{x})$ appears in theoretical statements (e.g.Eqs. (8) and (1)), it is not assumed to be known or used in training. Its primary role is in analysis: establishing properties like Bayes optimality, bounding calibration error, or characterizing over-/underconfidence. The actual training and evaluation rely exclusively on $\{(\boldsymbol{x}_{i},y_{i})\}$. Consequently, there is no contradiction between the theoretical need for $\boldsymbol{\eta}(\boldsymbol{x})$ in proofs and the final empirical solution (e.g. the surrogate risk minimization) that depends only on observable data.
3 Proposed Method
3.1 Focal Calibration Loss
Our observations in Fig. 2 suggest that focal loss often yields higher-entropy predictions, leading to underconfident outputs even when the true probability $\eta_{k}(\boldsymbol{x})$ is near 0 or 1 (i.e., for “easy” samples). Such underconfidence undermines calibration reliability. Although Charoenphakdee et al. (Charoenphakdee et al., 2021) proved that focal loss is classification-calibrated, it is not strictly proper. We address these calibration issues by proposing the Focal Calibration Loss (FCL), which appends a proper loss term to focal loss.
Figure 2: Binary classification ($K=2$): Optimal predicted probabilities $\hat{p}$ for ground-truth probabilities $q\in[0,1]$ using Cross-Entropy Loss, Focal Loss, and Focal Calibration Loss.
Definition 3.1 (Focal Calibration Loss (FCL)).
Let $f:\mathcal{X}\rightarrow\Delta^{K}$ map inputs $\boldsymbol{x}_{i}\in\mathcal{X}$ to predicted probabilities $\hat{\boldsymbol{p}}_{i}=f(\boldsymbol{x}_{i})=[\hat{p}_{i1},\hat{p}_{i2},\ldots,\hat{p}_{iK}]^{\top}\in\Delta^{K}$. Let $\boldsymbol{y}_{i}=[y_{i1},y_{i2},\ldots,y_{iK}]^{\top}\in\Delta^{K}$ be the one-hot encoded true label. We define the Focal Calibration Loss as
$\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}\;=\;\frac{1}{N}\sum_{i=1}^{N}\!\Bigl{[}\,\mathcal{L}_{\text{focal}}\bigl{(}f(\boldsymbol{x}_{i}),\boldsymbol{y}_{i}\bigr{)}\;+\;\lambda\,\mathcal{L}_{\text{calib}}\bigl{(}f(\boldsymbol{x}_{i}),\boldsymbol{y}_{i}\bigr{)}\Bigr{]},$ (10)
where $\gamma\geq 0$ and $\lambda\geq 0$ control the focal “focusing” and calibration trade-off, respectively. Specifically:
$\mathcal{L}_{\text{focal}}(f(\boldsymbol{x}_{i}),\boldsymbol{y}_{i})=-\sum_{k=1}^{K}\;y_{ik}\,\bigl{(}1-\hat{p}_{ik}\bigr{)}^{\gamma}\,\log\hat{p}_{ik},$ (11)
$\mathcal{L}_{\text{calib}}(f(\boldsymbol{x}_{i}),\boldsymbol{y}_{i})=\left\|f(\boldsymbol{x}_{i})-\boldsymbol{y}_{i}\right\|_{2}^{2}=\sum_{k=1}^{K}\bigl{(}\hat{p}_{ik}\;-\;y_{ik}\bigr{)}^{2}.$ (12)
In practice, $\boldsymbol{y}_{i}$ is the (one-hot) observed class label for $\boldsymbol{x}_{i}$. By contrast, $\boldsymbol{\eta}(\boldsymbol{x})$ in certain theoretical derivations is the true conditional probability distribution $P(y\mid\boldsymbol{x})$, which is unknown in real-world datasets. Eq. (10) thus provides a fully empirical training objective and we define $\sigma^{\gamma,\lambda}(q):=(1-q)^{\gamma}-\gamma\,q\log(q)\,(1-q)^{\gamma-1}-2\lambda\,q,$ whose strict monotonicity and unique root properties are established in Appendix A.6.
Why FCL Helps.
Adding the squared-error calibration term $\lambda\,\|\hat{\boldsymbol{p}}_{i}-\boldsymbol{y}_{i}\|_{2}^{2}$ directly rewards probability estimates that approach the one-hot label, thereby nudging predictions toward lower calibration error. Even though one-hot labels are a crude proxy for $\boldsymbol{\eta}(\boldsymbol{x})$, this step ensures that, over many samples, the model’s output distribution more closely aligns with empirical label frequencies.
4 Theoretical Evidence
4.1 Mitigating OC/UC Beyond Focal Loss
Theorem 4.1 (Reducing Over-/Underconfidence).
Let $f$ be a classifier trained by minimizing $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ with $\lambda>0$. Suppose $\boldsymbol{\eta}(\boldsymbol{x})$ is the (unknown) true conditional probability vector. Then for any $\boldsymbol{x}\in\mathcal{X}$,
$\bigl{|}\,\max_{k}\hat{p}_{k}(\boldsymbol{x})\;-\;\max_{k}\eta_{k}(\boldsymbol{x})\bigr{|}\;\leq\;\|\,\hat{\boldsymbol{p}}(\boldsymbol{x})\;-\;\boldsymbol{\eta}(\boldsymbol{x})\|_{2}.$ (13)
Although $\boldsymbol{\eta}(\boldsymbol{x})$ is not directly accessible, this theorem shows that, if the model predictions $\hat{\boldsymbol{p}}(\boldsymbol{x})$ approach $\boldsymbol{\eta}(\boldsymbol{x})$ in $\ell_{2}$-distance, the gap between their respective maximum components shrinks. As a result, reducing $\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\|_{2}$ (which is encouraged by $\mathcal{L}_{\text{calib}}$) inherently bounds over-/underconfidence. Formal details in Appendix A.2 and  A.3.
4.2 $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ is Classification-Calibrated
Next, we show that minimizing the focal calibration risk $\mathcal{R}^{\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}}$ achieves the Bayes-optimal classification boundary (Zhang, 2004). Although calibration term in Eq. (12) does not necessarily recover the exact posterior $\boldsymbol{\eta}(\boldsymbol{x})$, we prove it preserves the ranking of classes required for classification consistency (Bartlett et al., 2006; Tewari & Bartlett, 2007).
Definition 4.2 (Pointwise Conditional Risk).
Let $\mathcal{L}\colon\Delta^{K}\times\Delta^{K}\to\mathbb{R}$ be a surrogate loss, and let $\boldsymbol{e}_{y}\in\{0,1\}^{K}$ be the one-hot label for class $y$. For an input $\boldsymbol{x}\in\mathcal{X}$ with true class-posterior $\boldsymbol{\eta}(\boldsymbol{x})=[\eta_{1}(\boldsymbol{x}),\dots,\eta_{K}(\boldsymbol{x})]^{\top}$, and a predicted distribution $\boldsymbol{q}(\boldsymbol{x})\in\Delta^{K}$, the pointwise conditional risk is
$\mathcal{R}^{\mathcal{L}}\!\bigl{(}\boldsymbol{q}(\boldsymbol{x});\boldsymbol{\eta}(\boldsymbol{x})\bigr{)}\;=\;\sum_{y\in\mathcal{Y}}\eta_{y}(\boldsymbol{x})\,\mathcal{L}\bigl{(}\boldsymbol{q}(\boldsymbol{x}),\,\boldsymbol{e}_{y}\bigr{)}.$ (14)
Definition 4.3 (Classification-Calibrated (Bartlett et al., 2006; Tewari & Bartlett, 2007)).
A loss $\mathcal{L}$ is said to be classification-calibrated if its minimizer preserves the same $\arg\max$ ordering as the true posterior. Formally, let
$\boldsymbol{q}^{*}(\boldsymbol{x})=\arg\min_{\boldsymbol{q}\in\Delta^{K}}\mathcal{R}^{\mathcal{L}}\bigl{(}\boldsymbol{q}(\boldsymbol{x});\boldsymbol{\eta}(\boldsymbol{x})\bigr{)}.$ (15)
Then $\mathcal{L}$ is classification-calibrated if, for every $\boldsymbol{x}$, $\arg\max_{y}\,\bigl{[}q_{y}^{*}(\boldsymbol{x})\bigr{]}\;=\;\arg\max_{y}\,\bigl{[}\eta_{y}(\boldsymbol{x})\bigr{]}.$
Intuitively, if a surrogate loss is classification-calibrated, minimizing its pointwise risk yields a decision rule (via $\arg\max$) that matches the Bayes rule, hence achieving Bayes-optimal classification (Zhang, 2004). In Appendix A.5, we prove that $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ satisfies the strictly order-preserving property, ensuring $\arg\max_{y}\hat{p}_{y}(\boldsymbol{x})$ coincides with $\arg\max_{y}\eta_{y}(\boldsymbol{x})$. Consequently, $\mathcal{R}^{\mathcal{L}_{0\text{-}1}}\bigl{(}f^{\boldsymbol{q}^{\gamma,\lambda,*}}\bigr{)}\;=\;\mathcal{R}^{\mathcal{L}_{0\text{-}1}}\bigl{(}f^{\mathcal{L}_{0\text{-}1},*}\bigr{)},$ indicating that $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ recovers the Bayes-optimal decision boundary. Thus, by minimizing FCL, we obtain a classifier whose $\arg\max$ matches the optimal $\arg\max$ implied by $\boldsymbol{\eta}$, meeting the requirement for Def. 4.3.
4.3 Strictly Proper
Although classification calibration ensures correct $\arg\max$ decisions, many real-world tasks also require accurate probability estimates. A loss $\mathcal{L}$ is strictly proper if it is minimized only when $\hat{\boldsymbol{p}}(\boldsymbol{x})$ exactly matches the true probability $\boldsymbol{\eta}(\boldsymbol{x})$. While the true $\boldsymbol{\eta}(\boldsymbol{x})$ remains unknown, strictly proper losses incentivize unbiased probability estimation (Reid & Williamson, 2010; Williamson et al., 2016).
Definition 4.4 (Strictly Proper Loss).
$\mathcal{L}$ is strictly proper if, for any $\boldsymbol{y}\in\Delta^{K}$ and any $\hat{\boldsymbol{p}}\neq\boldsymbol{y}$, $\mathcal{L}(\boldsymbol{y},\boldsymbol{y})\;<\;\mathcal{L}(\hat{\boldsymbol{p}},\boldsymbol{y}).$
We show in Appendix A.7 that the calibration term $\|\hat{\boldsymbol{p}}-\boldsymbol{y}\|_{2}^{2}$ is strictly proper. We further prove in Appendix A.8 that FCL is strictly proper if $\lambda>0$. Intuitively, adding $\|\hat{\boldsymbol{p}}-\boldsymbol{y}\|_{2}^{2}$ penalizes deviations from the observed label distribution more severely as $\hat{\boldsymbol{p}}$ drifts away, forcing the model to converge to correct probability estimates in expectation.
4.4 Reduction of Post-Processing Gap
We finally show that FCL reduces the post-processing gap (Blasiok et al., 2024) relative to focal loss alone. Recall that post-processing gap measures how much a smooth recalibration function could further improve the model’s calibration:
Theorem 4.5.
For any predictor $f:\mathcal{X}\to[0,1]$ and any distribution $\mathcal{D}$ over $\mathcal{X}\times\{0,1\}$, the $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ yields a smaller (or equal) post-processing gap than focal loss alone:
$\text{pGap}_{\mathcal{D}}\!\bigl{(}\mathcal{L}_{\text{FCL}}\bigr{)}\;\leq\;\text{pGap}_{\mathcal{D}}\!\bigl{(}\mathcal{L}_{\text{focal}}\bigr{)}.$ (16)
This implies that by incorporating $\|\hat{\boldsymbol{p}}-\boldsymbol{y}\|_{2}^{2}$ into training, the model’s raw outputs require less recalibration after the fact. Additional details and proofs are in Appendix A.14.
5 Experimental Results
We evaluate the effectiveness of our proposed method on both image and text classification tasks. For image classification, we utilize datasets such as CIFAR-10/100 (Krizhevsky, 2009), Tiny-ImageNet (Deng et al., 2009), and CheXNet (Rajpurkar et al., 2017) to evaluate calibration of models including ResNet50 (He et al., 2016a), ResNet-100 (He et al., 2016b), Wide-ResNet-26-10 (Zagoruyko & Komodakis, 2016), and DenseNet-121 (Huang et al., 2017). For text classification, we employ the 20 Newsgroups dataset (Lang, 1995), web news dataset AG News (Zhang et al., 2015) and FinSen (Liang et al., 2024), training with global-pooling CNN (Lin et al., 2014). More, we include Out-of-Distribution (OoD) datasets to evaluate the robustness of the models: the SVHN dataset (Netzer et al., 2011), which consists of street view house numbers, and the CIFAR-10-C dataset (Hendrycks & Dietterich, 2019), a corrupted version of the CIFAR-10.
Baselines Among calibration-during-training methods, Cross Entropy with Weight Decay ($5\times 10^{-4}$)(Guo et al., 2017), MMCE (Kumar et al., 2018), Brier loss (Brier, 1950), Label Smoothing (LS-0.05) (Szegedy et al., 2016), sample-dependent focal loss (FLSD-53)(Lin et al., 2017), and Dual Focal Loss (Tao et al., 2023) as baselines. For post hoc calibration, we report the effect of temperature scaling(Guo et al., 2017) on top of these methods. For temperature scaling, we select the optimal temperature $T\in(0,10]$ (step size 0.1) that gives the lowest ECE on the validation set.
Experiment Setup Our experiment setup aligns with prior work and public code by Mukhoti et al. (Mukhoti et al., 2020a) implemented using PyTorch. CIFAR-10/100: Trained ResNet-50, ResNet-110, WideResNet, DenseNet for 350 epochs, with a learning rate set to 0.1 for the first 150 epochs, 0.01 for the following 100 epochs, and 0.001 for the remaining epochs. Tiny-ImageNet: Trained ResNet for 100 epochs, with a learning rate set to 0.1 for the first 40 epochs, 0.01 for the following 20 epochs, and 0.001 for the remaining epochs. 20 Newsgroups: Trained Global Pooling CNN for 50 epochs, with a learning rate set to 0.001, beta set from 0.9 to 0.999. We used Glove word embedding (glove.6B.100d) to train and evaluated by the best validated models. CheXNet: Trained DenseNet for 100 epochs with a ReduceLROnPlateau scheduler (patience = 5). For optimization, we used SGD with a weight decay of $5\times 10^{-4}$ and momentum of 0.9 for CIFAR-10/100 and Tiny-ImageNet, Adam optimizer with a learning rate of 0.001 for 20Newsgroup, and Adam optimizer with a learning rate of 0.0001 for CheXNet with $1\times 10^{-5}$ weight decay. The batch sizes are all set to 128 but 32 for CheXNet. All experiments were conducted on NVIDIA GeForce RTX 3060 Ti and NVIDIA GeForce RTX 4090, with random seeds set to 1. For temperature scaling results, the temperature parameter $T$ was optimized through a grid search over the range $T\in[0,0.1,0.2,\ldots,10]$ on the validation set, based on the best post-temperature-scaling ECE. The exact temperature parameter was also used for other calibration metrics.
5.1 Calibration Performance
5.1.1 Calibration Error
Tables 4, 5, 6, and 7 present the ECE, smCE, AdaECE, and Classwise-ECE metrics, respectively, before (pre) and after (post) temperature scaling for different methods across various models and datasets. Overall, FCL consistently achieves lower pre-T calibration error compared to other methods, indicating better inherent calibration. Post-T calibration error for FCL are also among the lowest, showcasing its effectiveness even after temperature scaling. FCL performs well across different architectures, particularly excelling with DenseNet-121 and ResNet-50/110. It demonstrates strong performance on CIFAR-10, CIFAR-100, and Tiny-ImageNet, with significant improvements in calibration. These results underscore the efficacy of FCL in enhancing model calibration, outperforming other methods like Brier Loss, MMCE, and Label Smoothing, Focal Loss in both pre- and post-temperature scaling scenarios.
Table 1: $\downarrow$ ECE (%) for various methods both before (pre) and after (post) applying temperature scaling (Bin = 15)  [^4] We follow the experimental methodology of (Mukhoti et al., 2020a), utilizing their publicly available code (Mukhoti et al., 2020b). For the Dual Focal Loss, we initially set $\gamma=5$ based on preliminary tuning reported by (Tao et al., 2023) to ensure a fair comparison. Since specific $\gamma$ values for other models and datasets were not provided, we performed cross-validation to determine the optimal $\gamma$ for each experiment. These optimized $\gamma$ values may influence the observed results by enhancing model calibration and performance. Optimal temperature scaling values determined by validation set (included in brackets) close to 1.0 indicate well-calibrated predictions requiring minimal adjustment..
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T
CIFAR-100 ResNet-50 18.02 2.60(2.2) 5.47 3.71(1.1) 15.05 3.41(1.9) 6.38 5.12(1.1) 5.54 2.28(1.1) 8.79 2.31(1.3) 3.71 2.14(1.1)
ResNet-110 19.29 4.75(2.3) 6.72 3.59(1.2) 18.84 4.52(2.3) 9.53 5.20(1.3) 11.02 3.88(1.3) 11.65 3.75(1.3) 4.19 3.23(1.1)
Wide-ResNet-26-10 15.17 2.82(2.1) 4.07 3.03(1.1) 13.57 3.89(2.0) 3.29 3.29(1.0) 2.42 2.16(1.1) 5.30 2.27(1.2) 2.22 2.22(1.0)
DenseNet-121 19.07 3.42(2.2) 4.28 2.22(1.1) 17.37 3.22(2.0) 8.63 4.82(1.1) 3.40 1.55(1.1) 6.69 1.65(1.2) 1.31 1.31(1.0)
CIFAR-10 ResNet-50 4.23 1.37(2.5) 1.83 0.96(1.1) 4.67 1.11(2.6) 4.23 1.87(0.9) 1.46 1.46(1.0) 1.32 1.32(1.0) 0.76 0.76(1.0)
ResNet-110 4.86 1.94(2.6) 2.50 1.36(1.2) 5.22 1.24(2.8) 3.74 1.22(0.9) 1.67 1.10(1.1) 1.48 1.27(1.1) 0.95 0.95(1.0)
Wide-ResNet-26-10 3.24 0.86(2.2) 1.07 1.07(1.0) 3.60 1.26(2.2) 4.66 1.27(0.8) 1.83 1.17(0.9) 3.35 0.94(0.8) 0.92 0.92(1.0)
DenseNet-121 4.70 1.54(2.4) 1.24 1.24(1.0) 4.97 1.71(2.4) 4.05 1.01(0.9) 1.73 1.22(0.9) 0.70 1.31(0.9) 0.66 0.66(1.0)
Tiny-ImageNet ResNet-50 16.03 5.23(1.5) 5.12 3.49(0.9) 13.15 4.77(1.3) 15.05 5.35(0.7) 1.61 1.61(1.0) 2.69 2.69(1.0) 10.22 1.35(0.8)
NLP 20 Newsgroups Global Pooling CNN 17.92 2.39(2.3) 15.48 6.78(2.1) 15.38 3.22(1.9) 4.79 2.54(1.1) 6.92 2.19(1.1) 18.45 3.89(2.3) 15.35 1.54(2.2)
NLP AG News Global Pooling CNN 4.82 0.67(2.0) 6.83 2.73(2.8) 1.87 1.79(1.1) 3.90 1.62(0.8) 6.70 0.87(3.0) 2.87 2.44(0.8) 0.99 0.34(1.1)
NLP FinSen Global Pooling CNN 0.98 0.44(2.5) 0.54 0.54(1.0) 1.23 0.92(0.5) 4.41 0.53(0.6) 0.43 0.53(0.9) 0.40 0.50(0.7) 0.51 0.31(1.3)
Table 2: $\downarrow$ Classification error (%) on test set across different methods.
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
CIFAR-100 ResNet-50 22.75 25.29 22.16 23.03 22.44 21.97 21.81
ResNet-110 22.93 24.79 22.78 22.98 22.04 21.73 21.02
Wide-ResNet-26-10 20.64 20.45 20.59 20.36 19.89 20.05 19.81
DenseNet-121 24.20 23.83 23.70 21.85 22.84 23.00 21.80
CIFAR-10 ResNet-50 4.87 5.17 5.14 4.72 4.94 5.70 4.56
ResNet-110 5.20 5.34 5.56 4.70 5.19 5.82 4.75
Wide-ResNet-26-10 3.80 4.46 4.22 4.12 4.34 4.12 4.16
DenseNet-121 5.11 4.82 5.36 4.79 5.46 5.39 4.59
Tiny-ImageNet ResNet-50 50.36 53.47 51.57 47.93 48.84 49.73 47.68
20 Newsgroups Global Pooling CNN 27.66 27.16 28.41 28.03 31.16 28.93 27.13
AG News Global Pooling CNN 8.86 8.53 8.21 8.33 8.84 8.28 8.13
FinSen Global Pooling CNN 1.13 1.26 1.42 1.13 1.22 1.09 1.22
Table 3: $\downarrow$ smCE (%) for various methods both before (pre) and after (post) applying temperature scaling  [^5] The smCE metrics is based on the approach presented by (Błasiok & Nakkiran, 2023), with implementation publicly on GitHub (Błasiok, Jarosław and Nakkiran, Preetum, 2024)..
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T
CIFAR-100 ResNet-50 14.95 2.63(2.2) 5.34 3.53(1.1) 13.40 3.12(1.9) 6.31 3.43(1.1) 5.56 2.37(1.1) 8.82 2.24(1.3) 3.68 2.29(1.1)
ResNet-110 15.05 3.84(2.3) 6.53 3.51(1.2) 14.86 3.51(2.3) 9.55 4.36(1.3) 10.98 3.60(1.3) 11.69 3.31(1.3) 4.16 3.08(1.1)
Wide-ResNet-26-10 12.97 2.85(2.1) 4.06 2.78(1.1) 12.27 3.85(2.0) 3.04 3.04(1.0) 2.34 2.20(1.1) 6.01 2.41(0.9) 2.16 2.16(1.0)
DenseNet-121 15.42 2.61(2.2) 3.86 1.93(1.1) 15.96 2.95(2.0) 3.85 3.68(1.1) 3.04 1.50(1.1) 4.07 1.72(0.9) 1.48 1.48(1.0)
CIFAR-10 ResNet-50 3.27 1.48(2.5) 1.83 1.27(1.1) 3.41 1.47(2.6) 3.31 1.92(0.9) 1.45 1.45(1.0) 1.36 1.36(1.0) 0.99 0.99(1.0)
ResNet-110 3.43 1.88(2.6) 2.49 1.63(1.2) 3.38 1.55(2.8) 2.87 1.52(0.9) 1.69 1.39(1.1) 1.48 1.35(1.1) 1.17 1.17(1.0)
Wide-ResNet-26-10 2.70 1.25(2.2) 1.55 1.55(1.0) 3.03 1.46(2.2) 3.75 1.47(0.8) 1.84 1.46(0.9) 3.31 1.20(0.8) 1.16 1.16(1.0)
DenseNet-121 15.42 2.61(2.4) 3.86 1.93(1.0) 15.96 2.95(2.4) 2.84 1.63(0.9) 3.04 1.50(0.9) 4.07 1.72(0.9) 1.48 1.48(1.0)
Tiny-ImageNet ResNet-50 15.80 5.02(1.5) 5.02 3.50(0.9) 13.02 4.65(1.3) 14.99 5.24(0.7) 1.46 1.46(1.0) 6.78 1.99(0.9) 10.23 1.39(0.8)
NLP 20 Newsgroups Global Pooling CNN 17.33 2.41(2.7) 15.13 1.85(2.2) 13.51 2.17(2.1) 4.77 7.40(0.8) 6.20 4.19(1.1) 10.95 3.54(1.8) 14.91 1.59(2.2)
NLP AG News Global Pooling CNN 4.56 0.80(2.0) 5.63 2.72(2.8) 1.78 1.75(1.1) 3.83 1.55(0.8) 6.07 0.75(3.0) 2.74 2.34(0.8) 1.12 0.60(1.1)
NLP FinSen Global Pooling CNN 0.88 0.66(2.5) 0.76 0.76(1.0) 1.30 0.91(0.5) 4.44 0.65(0.6) 0.68 0.68(0.9) 0.63 0.66(0.7) 0.72 0.65(1.3)
5.1.2 Reliability Diagram
The Fig. 9 for ResNet-110 trained on CIFAR-10 show the reliability diagram for different loss functions: Cross Entropy, FLSD-53, Dual Focal, and Focal Calibration. The Cross Entropy loss shows a significant gap between the expected and actual accuracy, especially in the higher confidence bins indicated poor calibration. Focal Calibration Loss achieves the best calibration among the methods, with the lowest ECE and a relatively low MCE. The expected and actual accuracies are closely aligned, indicating excellent calibration. Additionally, the lowest test error is demonstrating that FCL not only improves calibration but also maintains or improves classification performance as Fig. 8 for ResNet-50 trained on CIFAR-10 with epochs range from 100 to 300.
5.2 Classification Performance
5.2.1 Classification Error
The Tab. 2 presents the classification error (%) on the test set for various methods across different datasets and models. FCL consistently achieves the lowest classification error in most cases across all datasets and models, demonstrating its effectiveness in reducing test errors, especially in CIFAR-100 and News datasets.
5.3 Gradient-Based Localization
5.3.1 Grad-CAM Heatmaps on CheXNet
We evaluate the effects of Cross Entropy, Focal Loss, and Focal Calibration Loss on localization performance using Grad-CAM heatmaps, a popular visualization technique for identifying regions of interest in deep learning models (Selvaraju et al., 2017). Fig. 1 presents a comparison of heatmaps generated with each loss function applied to CheXNet (Rajpurkar et al., 2017), a deep learning model for chest X-ray classification. For Cross Entropy, the heatmaps show broader regions of activation, which suggest that the model has less certainty in localizing the anomalies. In contrast, focal loss demonstrates more concentrated regions of activation. This suggests that focal loss sharpens the model’s focus on challenging examples. However, while the activations are more focused, they may not always be aligned with precise boundaries, which could affect the confidence in anomaly localization. Focal Calibration Loss, on the other hand, provides the most well-defined and precise activation areas. The heatmaps generated show a balanced focus, particularly in difficult cases, while also improving probability calibration. This results in clearer boundaries and more accurate localization of anomalies lead to more confident and reliable predictions, making it especially suited for medical diagnosis tasks where precise localization and confidence in predictions are critical. Thus, FCL not only enhances model calibration but also refines anomaly localization, offering substantial benefits for applications in medical imaging where accurate and confident diagnoses are of utmost importance. CheXNet result shows in Appendix A.16.
5.4 Robustness on OoD Data Shift
We assess the robustness of models trained on in-distribution CIFAR-10 and evaluated on two distinct out-of-distribution (OoD) datasets: SVHN and CIFAR-10-C. Tab. 4 presents the performance of Area Under the Receiver Operating Characteristic curve (AUROC) under data shifts. Across most datasets, FCL consistently achieves the highest AUROC, outperforming other models in handling OoD data. This indicates that FCL enhances the model’s ability to distinguish between in-distribution and out-of-distribution samples effectively. This suggests that the FCL generalize well across different types of OoD shifts.
Table 4: Robustness on OoD Dataset Shift. $\uparrow$ AUROC (%) for shifting CIFAR-10 (in-distribution) to SVHN and CIFAR-10-C (OoD).
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T
CIFAR-10/SVHN ResNet-50 91.53 91.95 95.11 95.26 75.73 75.19 85.07 85.08 90.72 90.72 93.54 93.82 92.28 92.28
ResNet-110 77.56 76.11 94.49 94.67 75.01 73.40 82.94 83.03 90.11 90.18 95.26 95.69 92.60 92.60
Wide-ResNet-26-10 87.12 87.25 92.75 92.75 93.31 93.07 88.39 88.42 93.45 93.47 92.63 92.60 95.77 95.77
DenseNet-121 92.93 91.57 95.62 95.62 90.42 89.34 74.43 74.88 93.01 92.95 93.93 93.99 96.20 96.20
CIFAR-10/CIFAR-10-C ResNet-50 83.09 81.29 89.91 89.89 87.81 88.13 67.98 68.34 91.37 91.37 87.79 87.64 89.66 89.66
ResNet-110 74.89 71.87 77.11 76.58 73.57 72.62 65.26 65.58 78.46 77.90 83.35 82.54 89.88 89.88
Wide-ResNet-26-10 80.67 79.79 89.44 89.44 76.80 76.34 84.43 84.47 82.91 83.51 84.05 83.89 90.98 90.98
DenseNet-121 86.96 80.91 53.87 53.87 69.15 65.23 72.26 72.55 90.35 90.51 82.90 82.32 90.57 90.57
5.5 Discussion
Figure 3: Normalized test loss across training epochs for different loss functions on CIFAR-10 using ResNet-110. The comparison includes Cross Entropy, Focal Loss ($\gamma$=3.0), Dual Focal Loss ( $\gamma$=2.0), and Focal Calibration Loss ($\gamma$=3.0). Focal Calibration Loss with faster convergence and stability.
5.5.1 Faster convergence
In our experiments, we compared the performance of various loss functions on the CIFAR-10 dataset using a ResNet-110 architecture. Specifically, we evaluated CE, FL ($\gamma$=3.0), DFL ($\gamma$=2.0), and FCL ($\gamma$=3.0). The results, depicted in Figure 3, illustrate the normalized test loss across training epochs. During the initial epochs from 0 to 50, CE demonstrated a relatively stable decrease, while the FL showed higher variability with sharp peaks and troughs, indicating potential overfitting on hard examples. The FCL shows faster convergence, reaching a low normalized test loss more quickly than the other loss functions. This indicates that FCL might be more effective in stabilizing and minimizing the loss during the initial training phase. From epochs 50 to 150, all loss functions exhibited a reduction in variability, but the FCL and DFL consistently showed the lowest and most stable loss values with similar pattern after epoch 150.
5.5.2 Balance between $\gamma$ and $\lambda$
We compared the FCL and FL with different $\lambda$ values while keeping $\gamma=3.0$ constant by metrics including NLL, ECE, AdaECE, and Classwise-ECE both before and after temperature scaling as well as test error for each $\lambda$ in Fig. 6. NLL decreases initially as $\lambda$ increases from 0 to 0.5 but shows fluctuations and a slight increase beyond $\lambda=1.5$. For post-temperature scaling, NLL values are generally lower, indicating improved calibration. ECE decreases significantly when $\lambda$ is 0.5, indicating improved calibration. However, calibration error increases for higher $\lambda$ values, suggesting that too much regularization might degrade calibration performance as depicted in Fig. 6. The experiment shows that $\lambda=0.5$ provides the best calibration performance across all ECE metrics before and after temperature scaling. Adjusting $\gamma$ influences how much the model concentrates on hard examples—higher $\gamma$ increases this focus, which can improve calibration but may also lead to overfitting if excessive. Modifying $\lambda$ controls the weight of the calibration loss component; increasing $\lambda$ emphasizes probability calibration by penalizing deviations between predicted and true probabilities, but overly large values can reduce classification accuracy. Therefore, appropriate tuning of these hyperparameters is crucial to achieving a balance between calibration and classification. We provide table of $\gamma$ and $\lambda$ selection with best calibration error performance in Appendix A.10 and an ablation study of L1/L2 norm in Appendix A.15.
5.5.3 Post-processing gap Reduction
To empirically validate the theoretical reduction in post-processing gap as Theorem 4.5, we compare the ECE/smCE/AdaECE/Classwise-ECE of models trained with focal loss and FCL before and after temperature scaling. As shown in Tab. 4, Tab. 5, Tab. 6 and Tab.7, the focal loss model exhibits a significant reduction in ECE after temperature scaling, indicating a substantial post-processing gap. In contrast, the FCL model shows minimal change in calibration error after temperature scaling closing with 1.0, confirming that FCL reduces the post-processing gap.
Figure 4: Different hyperparameters trained on CIFAR-10 with ResNet-110. The plots show the performance of the FCL with varying calibration parameter $\lambda$ ($\gamma=3.0$) across different metrics: NLL, ECE, AdaECE, and Classwise-ECE. Each plot includes before (blue solid lines) and after (orange dashed lines) temperature scaling, alongside the Test Error (red dotted lines)  [^6] Our FCL reduces to the standard focal loss when $\lambda$ = 0..
6 Related Work
Improving the calibration of deep neural networks (DNNs) is critical for reliable uncertainty estimation. A common strategy involves post-hoc calibration methods, which adjust a pre-trained model’s output probabilities without altering its architecture or training procedure. Platt scaling (Platt et al., 1999), originally designed for binary classifiers, was extended to multi-class settings by applying a linear transformation to logits to improve calibration. More recent methods like Temperature Scaling (Guo et al., 2017) introduce a temperature parameter to adjust the predicted logits, effectively reducing the Expected Calibration Error (ECE) across various models. Histogram Binning (Zadrozny & Elkan, 2002) discretizes the probability space into bins to align confidence scores with empirical frequencies. While effective, these methods require held-out validation data and assume the test distribution matches the calibration data.
6.1 Calibration Strategies
Regularization-based approaches aim to improve calibration during training. Label Smoothing (Müller et al., 2019; Szegedy et al., 2016) mitigates overconfidence by blending true labels with a uniform distribution, softening hard targets to enhance generalization and uncertainty estimation. Additionally, auxiliary losses like the Maximum Mean Calibration Error (MMCE) (Kumar et al., 2018) explicitly minimize the difference between predicted and actual confidence levels. Bayesian techniques offer a principled way to quantify uncertainty. For instance, Deep Ensembles (Lakshminarayanan et al., 2017) aggregate predictions from multiple independently trained models to yield well-calibrated confidence estimates, though they are computationally intensive and less practical in resource-limited settings. Focal Loss (Lin et al., 2017), designed to handle class imbalance in object detection, also affects calibration by concentrating on hard-to-classify examples, which can reduce overconfidence but sometimes result in underconfident predictions (Charoenphakdee et al., 2021). Dual Focal Loss (DFL) (Tao et al., 2023) improves upon this by maximizing the margin between the true class logit and the next highest logit, balancing overconfidence and underconfidence to enhance calibration. Recently, smoothing-based methods have been introduced to provide continuous calibration measures. The Smooth Calibration Error (smCE) (Błasiok & Nakkiran, 2023) uses kernel-based smoothing to create a differentiable measure of calibration error, overcoming the limitations of discrete binning in traditional ECE calculations.
6.2 Calibration in Transformers
Calibration for transformer-based architectures, now prevalent in both vision and language domains, is increasingly essential for reliable deployment in high-stakes applications. In NLP, foundational transformer models such as BERT (Devlin, 2018), GPT (Brown, 2020), and T5 (Raffel et al., 2020) often lead the field in accuracy yet suffer from miscalibrated confidence (Desai & Durrett, 2020; Jiang et al., 2020), especially under out-of-distribution conditions (Mielke et al., 2022). Likewise, vision transformers (ViTs) exhibit calibration issues under distribution shifts (Ovadia et al., 2019), adversarial attacks (Shao et al., 2021), and low-data scenarios (Minderer et al., 2021). These limitations underscore the need for robust calibration methods to enhance reliability in domains such as autonomous driving (Feng et al., 2021) and medical imaging (Mehrtash et al., 2020).
7 Conclusions
In conclusion, traditional Focal Loss has been instrumental in reducing misclassification rates by focusing on hard-to-classify samples; however, it often falls short in providing well-calibrated predictions. By ensuring that the loss function is strictly proper, FCL delivers more reliable class-posterior probabilities, which directly contributes to improved model confidence and reliability across various applications even in unforeseen circumstances.
8 Impact Statement
Ensuring accurate and trustworthy confidence scores in deep neural networks is vital for deploying them in high-stakes domains such as healthcare, autonomous driving, and finance. By introducing a focal calibration loss that leverages an explicit Euclidean distance term, our approach not only addresses class imbalance but also reduces miscalibration, aligning model outputs with real-world outcome frequencies. This dual improvement in accuracy and reliability fosters greater transparency and confidence when models are used to assist critical decisions. Consequently, our method’s blend of focal weighting and calibration stands to significantly enhance model trustworthiness across a range of practical applications.
Appendix A Appendix
A.1 Evaluation Metrics
Formally, the ECE can be approximated by grouping predictions into $M$ bins and calculating the average confidence and accuracy for each bin:
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|$ (17)
where $N$ is the total number of samples, $M$ is the number of bins, $B_{m}$ is the set of indices of samples whose predicted probability falls into the $m$-th bin, $|B_{m}|$ is the number of samples in bin $B_{m}$, $\text{acc}(B_{m})$ is the accuracy of the samples in bin $B_{m}$, and $\text{conf}(B_{m})$ is the average confidence of the samples in bin $B_{m}$ defined in Eq. (18).
$\text{acc}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\mathbf{1}(\hat{y}_{i}=y_{i})\hfill,\quad\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p}_{i}$ (18)
$\begin{aligned} \text{MCE}=\max_{m\in{1,\ldots,M}}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|\end{aligned}\hfill$ (19)
$\begin{aligned} \text{AdaECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|\end{aligned}\hfill$ (20)
$\begin{aligned} \text{ClasswiseECE}=\frac{1}{K}\sum_{k=1}^{K}\sum_{m=1}^{M}\frac{|B_{km}|}{N_{k}}\left|\text{acc}(B_{km})-\text{conf}(B_{km})\right|\end{aligned}$ (21)
where $N_{k}$ is the number of samples in class $k$, $B_{km}$ is the set of indices of samples of class $k$ whose predicted probability falls into the $m$-th bin, $|B_{km}|$ is the number of samples in bin $B_{km}$, $\text{acc}(B_{km})$ is the accuracy of the samples in bin $B_{km}$, and $\text{conf}(B_{km})$ is the average confidence of the samples in bin $B_{km}$.
A.2 Proof of Thm. 4.1 Focal Calibration Loss mitigates overconfidence (OC) and underconfidence (UC)
In this appendix, we provide a mathematical proof that minimizing the Focal Calibration Loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ mitigates overconfidence (OC) and underconfidence (UC) in the predicted probabilities of a classifier.
Step 1: Bounding Overconfidence and Underconfidence
We aim to bound the difference between the maximum predicted probability and the maximum true class probability. For any $\boldsymbol{x}\in\mathcal{X}$, we have:
$\left|\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\right|\leq\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{\infty},$ (22)
where $\|\cdot\|_{\infty}$ denotes the infinity norm.
Proof.
Let $k^{*}=\arg\max_{k}\hat{p}_{k}(\boldsymbol{x})$ and $k^{\dagger}=\arg\max_{k}\eta_{k}(\boldsymbol{x})$. Then,
$\left|\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\right|=\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{\dagger}}(\boldsymbol{x})\right|.$ (23)
Using the triangle inequality, we have:
$\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{\dagger}}(\boldsymbol{x})\right|\leq\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{*}}(\boldsymbol{x})\right|+\left|\eta_{k^{*}}(\boldsymbol{x})-\eta_{k^{\dagger}}(\boldsymbol{x})\right|.$ (24)
Since $\eta_{k^{\dagger}}(\boldsymbol{x})\geq\eta_{k^{*}}(\boldsymbol{x})$ by the definition of $k^{\dagger}$ and $k^{*}$, it follows that:
$\left|\eta_{k^{*}}(\boldsymbol{x})-\eta_{k^{\dagger}}(\boldsymbol{x})\right|=\eta_{k^{\dagger}}(\boldsymbol{x})-\eta_{k^{*}}(\boldsymbol{x})\geq 0.$ (25)
Similarly, since $\hat{p}_{k^{*}}(\boldsymbol{x})\geq\hat{p}_{k^{\dagger}}(\boldsymbol{x})$, we have:
$\hat{p}_{k^{*}}(\boldsymbol{x})-\hat{p}_{k^{\dagger}}(\boldsymbol{x})\geq 0.$ (26)
Combining Equations (24) and (25), we get:
$\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{\dagger}}(\boldsymbol{x})\right|\leq\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{*}}(\boldsymbol{x})\right|+\left(\eta_{k^{\dagger}}(\boldsymbol{x})-\eta_{k^{*}}(\boldsymbol{x})\right).$ (27)
Since $\eta_{k^{\dagger}}(\boldsymbol{x})-\eta_{k^{*}}(\boldsymbol{x})\leq\max_{k}\left|\eta_{k}(\boldsymbol{x})-\eta_{k}(\boldsymbol{x})\right|=0$, the second term is zero. Therefore,
$\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{\dagger}}(\boldsymbol{x})\right|\leq\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{*}}(\boldsymbol{x})\right|.$ (28)
Thus,
$\left|\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\right|\leq\max_{k}\left|\hat{p}_{k}(\boldsymbol{x})-\eta_{k}(\boldsymbol{x})\right|=\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{\infty}.$ (29)
Step 2: Relating Infinity Norm to Euclidean Norm
We have the relationship between norms:
$\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{\infty}\leq\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{2}.$ (30)
This follows because the infinity norm is less than or equal to the Euclidean norm.
Step 3: Combining Steps 1 and 2
From inequalities (22) and (30), we obtain:
$\left|\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\right|\leq\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{2}.$ (31)
Step 4: Minimizing the Calibration Loss Reduces the Norm
The calibration loss for a single data point is:
$\mathcal{L}_{\text{calib}}(f(\boldsymbol{x}),\boldsymbol{y})=\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{y}\right\|_{2}^{2}.$ (32)
Since $\boldsymbol{y}$ is a one-hot encoding sampled from the distribution $\boldsymbol{\eta}(\boldsymbol{x})$, we have:
$\mathbb{E}_{\boldsymbol{y}\sim\boldsymbol{\eta}(\boldsymbol{x})}\left[\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{y}\right\|_{2}^{2}\right]=\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{2}^{2}+\sum_{k=1}^{K}\eta_{k}(\boldsymbol{x})(1-\eta_{k}(\boldsymbol{x})).$ (33)
The term $\sum_{k=1}^{K}\eta_{k}(\boldsymbol{x})(1-\eta_{k}(\boldsymbol{x}))$ is independent of $\hat{\boldsymbol{p}}(\boldsymbol{x})$ and represents the intrinsic uncertainty of the true distribution. Therefore, minimizing $\mathcal{L}_{\text{calib}}$ effectively minimizes $\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{2}^{2}$.
Step 5: Impact on Overconfidence and Underconfidence
By minimizing $\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{2}^{2}$, we directly reduce the bound in Equation (31) on $\left|\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\right|$. This means that the difference between the highest predicted probability and the highest true probability decreases, mitigating both overconfidence and underconfidence.
Minimizing the Focal Calibration Loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ includes minimizing the calibration loss component $\mathcal{L}_{\text{calib}}$, which reduces the Euclidean distance between the predicted probabilities $\hat{\boldsymbol{p}}(\boldsymbol{x})$ and the true class-posterior probabilities $\boldsymbol{\eta}(\boldsymbol{x})$. This, in turn, reduces the bound on the difference between the maximum predicted probability and the maximum true probability, thereby mitigating overconfidence and underconfidence.
A.3 Proof that Minimizing the Focal Calibration Loss Recovers the True Class-Posterior Probabilities
Proof.
Let $\mathcal{X}$ be the input space and $\mathcal{Y}=\{1,2,\dots,K\}$ the set of $K$ classes. For a given input $\boldsymbol{x}\in\mathcal{X}$, let $\boldsymbol{\eta}=[\eta_{1},\eta_{2},\dots,\eta_{K}]^{\top}$ denote the true class-posterior probabilities, where $\eta_{i}=\mathbb{P}(y=i\mid\boldsymbol{x})$ and $\sum_{i=1}^{K}\eta_{i}=1$; $\boldsymbol{q}=[q_{1},q_{2},\dots,q_{K}]^{\top}$ be the predicted probability vector, where $q_{i}\in(0,1)$ and $\sum_{i=1}^{K}q_{i}=1$. We aim to minimize the Focal Calibration Loss (FCL) and Our goal is to show that the minimizer $\boldsymbol{q}^{*}$ of $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ satisfies $q_{i}^{*}=\eta_{i}$ for all $i$.
$\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})=-\sum_{i=1}^{K}\eta_{i}(1-q_{i})^{\gamma}\log q_{i}+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2},$ (34)
subject to $q_{i}>0$ and $\sum_{i=1}^{K}q_{i}=1$.
Step 1: Convexity of the Loss Function
Before proceeding, we establish that $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ is strictly convex in $\boldsymbol{q}$ over the domain $q_{i}\in(0,1)$.
  • Calibration Loss Term: The term $\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2}$ is strictly convex in $\boldsymbol{q}$ because it is a sum of squared terms.
  • Focal Loss Term: The term $\sum_{i=1}^{K}\eta_{i}(1-q_{i})^{\gamma}\log q_{i}$ is convex in $q_{i}$ for $q_{i}\in(0,1)$. The function $\phi(q_{i})=-(1-q_{i})^{\gamma}\log q_{i}$ is convex in $q_{i}\in(0,1)$ for $\gamma\geq 0$. Since $\eta_{i}\geq 0$, and a non-negative weighted sum of convex functions is convex, the focal loss term is convex.
The sum of a strictly convex function (calibration loss) and a convex function (focal loss) is strictly convex (See Thm. A.1 in Appendix A.4). Therefore, $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ is strictly convex in $\boldsymbol{q}$.
Step 2: Existence and Uniqueness of the Minimizer
Because $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ is strictly convex and the feasible set defined by $q_{i}>0$ and $\sum_{i=1}^{K}q_{i}=1$ is convex and closed, there exists a unique global minimizer $\boldsymbol{q}^{*}$.
Step 3: Setting Up the Lagrangian
To find the minimizer, we set up the Lagrangian function incorporating the equality constraint:
$\mathcal{L}(\boldsymbol{q},\mu)=-\sum_{i=1}^{K}\eta_{i}(1-q_{i})^{\gamma}\log q_{i}+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2}+\mu\left(\sum_{i=1}^{K}q_{i}-1\right).$ (35)
We do not need to include inequality constraints $q_{i}>0$ explicitly because $q_{i}\in(0,1)$ due to the logarithm and the domain of $\boldsymbol{q}$.
Step 4: Deriving the KKT Conditions
The Karush-Kuhn-Tucker (KKT) conditions for optimality are:
  1. 1.
    Stationarity: For each $i$, $\frac{\partial\mathcal{L}}{\partial q_{i}}=0.$
  2. 2.
    Primal Feasibility: $\sum_{i=1}^{K}q_{i}=1.$
  3. 3.
    Dual Feasibility: Not applicable here since there are no inequality constraints in $q_{i}$ after considering the domain.
  4. 4.
    Complementary Slackness: Not applicable since we don’t have inequality constraints with slack variables.
Compute the partial derivative:
$\frac{\partial\mathcal{L}}{\partial q_{i}}=\eta_{i}\left[\gamma(1-q_{i})^{\gamma-1}\log q_{i}+\frac{(1-q_{i})^{\gamma}}{q_{i}}\right]+2\lambda(q_{i}-\eta_{i})+\mu=0.$ (36)
Step 5: Analyzing the Stationarity Condition
Let’s denote:
$S_{i}(q_{i})=\eta_{i}\left[\gamma(1-q_{i})^{\gamma-1}\log q_{i}+\frac{(1-q_{i})^{\gamma}}{q_{i}}\right]+2\lambda(q_{i}-\eta_{i}).$ (37)
[…]
$S_{i}(q_{i})=S_{j}(q_{j})\quad\forall i,j.$ (38)
Step 6: Showing that $q_{i}^{*}=\eta_{i}$ satisfies the conditions
We need evaluate $S_{i}(q_{i})$ at $q_{i}=\eta_{i}$:
1. When $\eta_{i}>0$ and $\eta_{i}\in(0,1)$: Compute $\log\eta_{i}$, $(1-\eta_{i})$, and $\gamma$. Substitute $q_{i}=\eta_{i}$ into $S_{i}(q_{i})$:
$S_{i}(\eta_{i})$ $=\eta_{i}\left[\gamma(1-\eta_{i})^{\gamma-1}\log\eta_{i}+\frac{(1-\eta_{i})^{\gamma}}{\eta_{i}}\right]+2\lambda(\eta_{i}-\eta_{i})$ (39)
$=\eta_{i}\left[\gamma(1-\eta_{i})^{\gamma-1}\log\eta_{i}+\frac{(1-\eta_{i})^{\gamma}}{\eta_{i}}\right].$ (40)
2. Observation: The term $S_{i}(\eta_{i})$ depends only on $\eta_{i}$ and is constant for each $i$. Therefore, $S_{i}(\eta_{i})+\mu=0$ holds for each $i$ with the same $\mu$.
3. Equality Across Classes: Since $S_{i}(\eta_{i})+\mu=0$ for all $i$, and $S_{i}(\eta_{i})$ is a constant for each $i$, the only way this can hold is if $\eta_{i}=\eta_{j}$ or the terms are adjusted by $\lambda$.
However, this suggests that $q_{i}=\eta_{i}$ satisfies the stationarity condition for some $\mu$, but we need to ensure that this solution indeed minimizes the loss function.
Step 7: Uniqueness and Minimization
Since $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ is strictly convex, any stationary point is the unique global minimizer. Let’s evaluate $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ at $\boldsymbol{q}=\boldsymbol{\eta}$:
1. Calibration Loss term is minimized to 0: $\lambda\sum_{i=1}^{K}(\eta_{i}-\eta_{i})^{2}=0.$ 2. Focal Loss term $\sum_{i=1}^{K}\eta_{i}(1-\eta_{i})^{\gamma}\log\eta_{i}$ is minimized when $q_{i}=\eta_{i}$ because the negative log-likelihood is minimized when the predicted probabilities match the true probabilities.
Since $\boldsymbol{q}=\boldsymbol{\eta}$ yields the minimal possible value of $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$, and the loss function is strictly convex, $\boldsymbol{q}^{*}=\boldsymbol{\eta}$ is the unique global minimizer.
Step 8: Handling the case when $\eta_{i}=0$ or $\eta_{i}=1$
[…]
In both cases, $q_{i}^{*}=\eta_{i}$ minimizes the loss.
Therefore, minimizing the Focal Calibration Loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ over the probability simplex recovers the true class-posterior probabilities $\boldsymbol{q}^{*}=\boldsymbol{\eta}$.
A.4 Convexity of the Sum of Strictly Convex and Convex Functions
Theorem A.1.
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a strictly convex function, and let $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a convex function. Then the function $h=f+g$ is strictly convex on $\mathbb{R}^{n}$.
Proof.
Let $x,y\in\mathbb{R}^{n}$ with $x\neq y$, and let $\theta\in(0,1)$.
Since $f$ is strictly convex, we have:
$f\big{(}\theta x+(1-\theta)y\big{)}<\theta f(x)+(1-\theta)f(y).$ (41)
Similarly, since $g$ is convex, we have:
$g\big{(}\theta x+(1-\theta)y\big{)}\leq\theta g(x)+(1-\theta)g(y).$ (42)
Adding the two inequalities, we obtain:
$\begin{split}h\big{(}\theta x+(1-\theta)y\big{)}&=f\big{(}\theta x+(1-\theta)y\big{)}+g\big{(}\theta x+(1-\theta)y\big{)}\\ &<\theta f(x)+(1-\theta)f(y)+\theta g(x)+(1-\theta)g(y)\\ &=\theta\big{(}f(x)+g(x)\big{)}+(1-\theta)\big{(}f(y)+g(y)\big{)}\\ &=\theta h(x)+(1-\theta)h(y).\end{split}$ (43)
This strict inequality shows that $h$ is strictly convex on $\mathbb{R}^{n}$. ∎
A.5 Proof that the Focal Calibration Loss Satisfies the Strictly Order-Preserving Property
Theorem A.2.
The Focal Calibration Loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ is classification-calibrated; that is, it satisfies the Strictly Order-Preserving Property.
Proof.
From Def. LABEL:def:Strictly_Order-Preserving_Property, we aim to show that the Focal Calibration Loss (FCL) satisfies:
$q_{i}^{*}<q_{j}^{*}\Rightarrow\eta_{i}<\eta_{j},$ (44)
where $\boldsymbol{q}^{*}=[q_{1}^{*},q_{2}^{*},\dots,q_{K}^{*}]^{\top}$ is the minimizer of the pointwise conditional risk $\mathcal{R}^{\mathcal{L}_{\text{FCL}}}(\boldsymbol{q})$.
Step 1: Define the Focal Calibration Loss (FCL)
The Focal Calibration Loss for a single sample $\boldsymbol{x}$ with true label $y\in\{1,2,\ldots,K\}$ is defined as:
$\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q},y)=-(1-q_{y})^{\gamma}\log q_{y}+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2},$ (45)
where:
  • $\boldsymbol{q}=[q_{1},q_{2},\dots,q_{K}]^{\top}$ is the predicted probability distribution over the $K$ classes.
  • $\eta_{i}=\mathbb{P}(y=i\mid\boldsymbol{x})$ is the true class-posterior probability for class $i$.
  • $\gamma\geq 0$ is the focusing parameter.
  • $\lambda>0$ is the calibration coefficient.
Step 2: Define the Pointwise Conditional Risk
The pointwise conditional risk $\mathcal{R}^{\mathcal{L}}(\boldsymbol{q})$ for the loss function $\mathcal{L}$ is defined as:
$\mathcal{R}^{\mathcal{L}}(\boldsymbol{q})=\mathbb{E}_{y\mid\boldsymbol{x}}[\mathcal{L}(\boldsymbol{q},y)]=\sum_{i=1}^{K}\eta_{i}\mathcal{L}(\boldsymbol{q},i).$ (46)
For FCL, this becomes:
$\mathcal{R}^{\mathcal{L}_{\text{FCL}}}(\boldsymbol{q})=\sum_{i=1}^{K}\eta_{i}\left[-(1-q_{i})^{\gamma}\log q_{i}\right]+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2}.$ (47)
Step 3: Find the Minimizer of the Pointwise Conditional Risk
Our objective is to find the minimizer $\boldsymbol{q}^{*}$ of $\mathcal{R}^{\mathcal{L}_{\text{FCL}}}(\boldsymbol{q})$:
$\boldsymbol{q}^{*}=\arg\min_{\boldsymbol{q}}\left\{\mathcal{R}^{\mathcal{L}_{\text{FCL}}}(\boldsymbol{q})\mid\sum_{i=1}^{K}q_{i}=1,\quad q_{i}>0\quad\forall i\right\}.$ (48)
Step 4: Set Up the Lagrangian
Introduce a Lagrange multiplier $\mu$ for the equality constraint $\sum_{i=1}^{K}q_{i}=1$:
$\mathcal{L}(\boldsymbol{q},\mu)=\sum_{i=1}^{K}\eta_{i}\left[-(1-q_{i})^{\gamma}\log q_{i}\right]+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2}+\mu\left(\sum_{i=1}^{K}q_{i}-1\right).$ (49)
Step 5: Compute the Partial Derivatives
For each $i$, compute the partial derivative of $\mathcal{L}(\boldsymbol{q},\mu)$ with respect to $q_{i}$:
$\frac{\partial\mathcal{L}}{\partial q_{i}}=\eta_{i}\left[\gamma(1-q_{i})^{\gamma-1}\log q_{i}+\frac{(1-q_{i})^{\gamma}}{q_{i}}\right]+2\lambda(q_{i}-\eta_{i})+\mu.$ (50)
Set the derivative to zero at the minimizer $\boldsymbol{q}^{*}$:
$\frac{\partial\mathcal{L}}{\partial q_{i}}\bigg{|}_{\boldsymbol{q}=\boldsymbol{q}^{*}}=0.$ (51)
Thus:
$\eta_{i}\left[\gamma(1-q_{i}^{*})^{\gamma-1}\log q_{i}^{*}+\frac{(1-q_{i}^{*})^{\gamma}}{q_{i}^{*}}\right]+2\lambda(q_{i}^{*}-\eta_{i})+\mu=0.$ (52)
Step 6: Analyze the Stationarity Condition
From Equation (52), we observe that for each $i$:
$\eta_{i}\Phi(q_{i}^{*})+2\lambda(q_{i}^{*}-\eta_{i})+\mu=0,$ (53)
where:
$\Phi(q_{i}^{*})=\gamma(1-q_{i}^{*})^{\gamma-1}\log q_{i}^{*}+\frac{(1-q_{i}^{*})^{\gamma}}{q_{i}^{*}}.$ (54)
Since $\mu$ is the same for all $i$, we can write:
$\eta_{i}\Phi(q_{i}^{*})+2\lambda(q_{i}^{*}-\eta_{i})=-\mu\quad\forall i.$ (55)
Therefore, for any $i,j$:
$\eta_{i}\Phi(q_{i}^{*})+2\lambda(q_{i}^{*}-\eta_{i})=\eta_{j}\Phi(q_{j}^{*})+2\lambda(q_{j}^{*}-\eta_{j}).$ (56)
Step 7: Show the Strictly Order-Preserving Property
Suppose $q_{i}^{*}<q_{j}^{*}$. We need to show that $\eta_{i}<\eta_{j}$.
Assume, for contradiction, that $\eta_{i}\geq\eta_{j}$.
Case 1: $\eta_{i}=\eta_{j}$
From Equation (56), we have:
$\eta_{i}\left[\Phi(q_{i}^{*})-\Phi(q_{j}^{*})\right]+2\lambda(q_{i}^{*}-q_{j}^{*})=0.$ (57)
Since $q_{i}^{*}<q_{j}^{*}$ and $\eta_{i}>0$, we analyze the sign of $\Phi(q)$.
Observation:
  • The function $\Phi(q)$ is decreasing in $q$ over $q\in(0,1)$.
  • Therefore, $\Phi(q_{i}^{*})>\Phi(q_{j}^{*})$ since $q_{i}^{*}<q_{j}^{*}$.
  • $q_{i}^{*}-q_{j}^{*}<0$.
Thus, the left-hand side (LHS) is positive, which contradicts the equality to zero.
Case 2: $\eta_{i}>\eta_{j}$
Similarly, from Equation (56):
$\left[\eta_{i}\Phi(q_{i}^{*})-\eta_{j}\Phi(q_{j}^{*})\right]+2\lambda\left(q_{i}^{*}-\eta_{i}-q_{j}^{*}+\eta_{j}\right)=0.$ (58)
Since $q_{i}^{*}<q_{j}^{*}$ and $\eta_{i}>\eta_{j}$, we have:
  • $q_{i}^{*}-\eta_{i}<q_{j}^{*}-\eta_{j}$ because $q_{i}^{*}<q_{j}^{*}$ and $\eta_{i}>\eta_{j}$.
  • $\eta_{i}\Phi(q_{i}^{*})>\eta_{j}\Phi(q_{j}^{*})$ since $\eta_{i}>\eta_{j}$ and $\Phi(q_{i}^{*})>\Phi(q_{j}^{*})$.
Thus, the LHS is positive, leading to a contradiction.
Step 8: Conclusion
Our assumption that $\eta_{i}\geq\eta_{j}$ must be false. Therefore, if $q_{i}^{*}<q_{j}^{*}$, it must be that $\eta_{i}<\eta_{j}$.
Hence, the Focal Calibration Loss satisfies the Strictly Order-Preserving Property, and thus it is classification-calibrated. ∎
A.6 Properties of $\sigma^{\gamma,\lambda}$
Lemma A.3.
Let $\sigma^{\gamma,\lambda}(q)$ be defined as:
$\sigma^{\gamma,\lambda}(q)=(1-q)^{\gamma}-\gamma q\log q\,(1-q)^{\gamma-1}-2\lambda q,$ (59)
where $q\in(0,1)$, $\gamma\geq 0$, and $\lambda\geq 0$. Then, $\sigma^{\gamma,\lambda}(q)$ has the following properties:
  1. 1.
    $\sigma^{\gamma,\lambda}(q)$ is continuous and differentiable on $q\in(0,1)$.
  2. 2.
    As $q\to 0^{+}$, $\sigma^{\gamma,\lambda}(q)\to 1$.
  3. 3.
    As $q\to 1^{-}$, $\sigma^{\gamma,\lambda}(q)\to-2\lambda$.
  4. 4.
    $\sigma^{\gamma,\lambda}(q)$ is strictly decreasing on $q\in(0,1)$.
  5. 5.
    There exists a unique $q^{*}\in(0,1)$ such that $\sigma^{\gamma,\lambda}(q^{*})=0$.
Proof.
1. Continuity and Differentiability.
Each term in (59) is continuous and differentiable on $q\in(0,1)$:
  • $(1-q)^{\gamma}$ is continuous and differentiable for $q\in(0,1)$ and $\gamma\geq 0$.
  • $\gamma q\log q\,(1-q)^{\gamma-1}$ is continuous and differentiable since $\log q$ is continuous on $q\in(0,1)$.
  • $2\lambda q$ is a linear function and thus continuous and differentiable.
Therefore, $\sigma^{\gamma,\lambda}(q)$ is continuous and differentiable on $q\in(0,1)$.
2. Limit as $q\to 0^{+}$.
As $q\to 0^{+}$, we have $(1-q)^{\gamma}\to 1$, $q\log q\to 0$, and $2\lambda q\to 0$.
$\sigma^{\gamma,\lambda}(q)\to 1-0+0=1.$ (60)
3. Limit as $q\to 1^{-}$.
As $q\to 1^{-}$, $(1-q)^{\gamma}\to 0$, $\gamma q\log q\,(1-q)^{\gamma-1}\to 0$, and $2\lambda q\to-2\lambda$. Therefore:
$\sigma^{\gamma,\lambda}(q)\to 0+0-2\lambda=-2\lambda.$ (61)
4. Strict Monotonic Decrease.
Compute the first derivative $\sigma^{\prime\gamma,\lambda}(q)$:
$\begin{split}\sigma^{\prime\gamma,\lambda}(q)&=\frac{d}{dq}\left[(1-q)^{\gamma}-\gamma q\log q\,(1-q)^{\gamma-1}-2\lambda q\right]\\ &=-\gamma(1-q)^{\gamma-1}-\gamma\log q\,(1-q)^{\gamma-1}+\gamma q\frac{1}{q}(1-q)^{\gamma-1}\\ &\quad+\gamma(\gamma-1)q\log q\,(1-q)^{\gamma-2}-2\lambda\\ &=-\gamma(1-q)^{\gamma-1}-\gamma\log q\,(1-q)^{\gamma-1}+\gamma(1-q)^{\gamma-1}\\ &\quad+\gamma(\gamma-1)q\log q\,(1-q)^{\gamma-2}-2\lambda\\ &=\gamma\left[-(1-q)^{\gamma-1}+(1-q)^{\gamma-1}\right]-\gamma\log q\,(1-q)^{\gamma-1}\\ &\quad+\gamma(\gamma-1)q\log q\,(1-q)^{\gamma-2}-2\lambda\\ &=-\gamma\log q\,(1-q)^{\gamma-1}+\gamma(\gamma-1)q\log q\,(1-q)^{\gamma-2}-2\lambda.\end{split}$ (62)
Since $\log q<0$ for $q\in(0,1)$ and $(1-q)^{\gamma-1}>0$, the first term is positive. The second term is negative or zero depending on $\gamma$:
  • If $\gamma\geq 1$, then $(\gamma-1)\geq 0$, and since $\log q<0$, the second term is negative.
  • If $0\leq\gamma<1$, then $(\gamma-1)<0$, but $\log q<0$, so the product is positive; however, $q(1-q)^{\gamma-2}$ remains finite and positive.
Overall, $\sigma^{\prime\gamma,\lambda}(q)<0$ for all $q\in(0,1)$, implying that $\sigma^{\gamma,\lambda}(q)$ is strictly decreasing on $q\in(0,1)$.
5. Existence of a Unique Root.
Since $\sigma^{\gamma,\lambda}(q)$ is continuous on $[0,1]$ and strictly decreasing from $\sigma^{\gamma,\lambda}(0^{+})=1$ to $\sigma^{\gamma,\lambda}(1^{-})=-2\lambda$, by the Intermediate Value Theorem, there exists a unique $q^{*}\in(0,1)$ such that $\sigma^{\gamma,\lambda}(q^{*})=0$.
$\blacksquare$
Corollary A.4.
The equation $\sigma^{\gamma,\lambda}(q)=0$ has exactly one solution $q^{*}\in(0,1)$.
Understanding the properties of $\sigma^{\gamma,\lambda}(q)$ is crucial in analyzing the behavior of the Focal Calibration Loss. The function’s monotonic decrease ensures a unique mapping between the predicted probabilities $q$ and the true class probabilities $\eta$, aiding in establishing the consistency and calibration of the loss function.
Figure 5: Plot of $\sigma^{\gamma,\lambda}(q)$ vs. $q$ for different $\gamma$ and $\lambda$
The analysis of $\sigma^{\gamma,\lambda}(q)$ confirms that the function is strictly decreasing and crosses zero exactly once in the interval $q\in(0,1)$ in Fig. 5. This behavior is essential for ensuring that the Focal Calibration Loss provides a unique and consistent mapping between predicted probabilities and true class probabilities, contributing to the effectiveness of the loss function in classification tasks.
A.7 Proof: The calibration loss defined as the mean squared $L_{2}$ norm is a strictly proper scoring rule for both binary and multiclass classification.
Proof.
In the proof, $Y\sim P$ denotes a random variable $Y$ that follows the probability distribution $P$. The predicted probability for class $k$ is $q_{k}$, and the true probability for class $k$ is $p_{k}$. The expectation $\mathbb{E}_{Y\sim P}[S(Q,Y)]$ represents the expected value of the scoring function $S(Q,Y)$ over the distribution $P$. The calibration loss $\mathcal{L}_{\text{calib}}$ is defined as the mean squared $L_{2}$ norm, and it measures the squared differences between predicted probabilities $\hat{p}_{i}$ and actual binary outcomes $y_{i}$ for binary classification, or between predicted probability vectors $\hat{\boldsymbol{p}}_{i}$ and true outcome vectors $\boldsymbol{y}_{i}$ for multiclass classification. The derivatives $\frac{\partial}{\partial q_{j}}$ and $\frac{\partial^{2}}{\partial q_{j}^{2}}$ denote the first and second partial derivatives w.r.t. the predicted probability $q_{j}$, used to find and verify the maxima of the expected score.
The $L_{2}$ norm and the Brier score are closely related. For binary classification, the squared $L_{2}$ norm of the difference between predicted probabilities $\hat{p}_{i}$ and actual binary outcomes $y_{i}$ is equivalent to the Brier score:
$\frac{1}{N}\sum_{i=1}^{N}(\hat{p}_{i}-y_{i})^{2}$ (63)
For multiclass classification, the mean squared $L_{2}$ norm is:
$\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}(\hat{p}_{ik}-y_{ik})^{2}$ (64)
This is the Brier score for multiclass classification.
1) Binary Classification:
$\mathbb{E}_{Y\sim P}[S(Q,Y)]=-P(Y=1)(q-1)^{2}+P(Y=0)q^{2}$ (65)
Let $p$ be the true probability of the event $Y=1$. Then, the expected score is:
$\mathbb{E}_{Y\sim P}[S(Q,Y)]=-p(q-1)^{2}-(1-p)q^{2}$ (66)
To find the maxima, take the derivative w.r.t. $q$ and set it to zero:
$\begin{split}\frac{\partial}{\partial q}\mathbb{E}_{Y\sim P}[S(Q,Y)]&=-2p(q-1)-2(1-p)q\\ &=-2pq+2p-2q+2pq\\ &=2p-2q\Rightarrow 0=2p-2q\Rightarrow p=q\end{split}$ (67)
We need to check the second derivative to see if it is a maximum (for the properness condition) and if it is the only maximizer (for the strictness condition):
$\frac{\partial^{2}}{\partial q^{2}}\mathbb{E}_{Y\sim P}[S(Q,Y)]=-2<0$ (68)
The second derivative is always negative which means that the function is concave and the maximum is unique. Therefore, $p=q$ is the only maximizer and the Brier scoring rule for binary classification is strictly proper.
2) Multiclass Classification:
$\mathbb{E}_{Y\sim P}[S(Q,Y)]$ $=\sum_{k}P(Y=k)\left[-\sum_{i}(q_{i}-y_{i})^{2}\right]$ (69)
$=\sum_{k}P(Y=k)\left[-(q_{k}-1)^{2}-\sum_{i\neq k}q_{i}^{2}\right]$
$=\sum_{k}P(Y=k)\left[-(q_{k}-1)^{2}+q_{k}^{2}-\sum_{i}q_{i}^{2}\right]$
$=\sum_{k}P(Y=k)\left[-q_{k}^{2}-1+2q_{k}+q_{k}^{2}-\sum_{i}q_{i}^{2}\right]$
$=\sum_{k}P(Y=k)\left[2q_{k}-1-\sum_{i}q_{i}^{2}\right]$
$=\sum_{k}P(Y=k)(2q_{k}-1)-\sum_{i}q_{i}^{2}$
This expression can be expressed as follows by replacing $q_{iK}$ with $1-\sum_{i\neq K}q_{i}$:
$\mathbb{E}_{Y\sim P}[S(Q,Y)]=p_{1}(2q_{1}-1)-q_{1}^{2}+p_{2}(2q_{2}-1)-q_{2}^{2}+\ldots+p_{K}(2q_{K}-1)-q_{K}^{2}$ (70)
Taking the derivative w.r.t. $q_{j}$ and setting it to zero, we obtain:
$\frac{\partial}{\partial q_{j}}\mathbb{E}_{Y\sim P}[S(Q,Y)]$ $=2p_{j}-2q_{j}-2p_{K}+2\left(1-\sum_{i\neq K}q_{i}\right)$ (71)
$=2p_{j}-2q_{j}-2p_{K}+2q_{K}$
$=(p_{j}-q_{j})+(q_{K}-p_{K})=0$
$(p_{j}-q_{j})=(p_{K}-q_{K})$
We know that $\sum_{i}p_{i}=1$ and $\sum_{i}q_{i}=1$, therefore:
$p_{1}-q_{1}$ $=p_{2}-q_{2}=\ldots=p_{K}-q_{K}=\lambda$ (72)
$\sum_{i}(p_{i}-q_{i})=K\cdot\lambda=0$
$\Rightarrow\lambda=0\quad\text{since}\quad K\neq 0$
$\Rightarrow p_{i}=q_{i}\quad\text{for all}\quad i=1,\ldots,K$
Now, we need to check the second derivative to see if it is a maximum (for the properness condition) and if it is the only maximizer for the strictness condition:
$\frac{\partial^{2}}{\partial q_{j}^{2}}\mathbb{E}_{Y\sim P}[S(Q,Y)]=-2-2=-4<0$ (73)
The second derivative is always negative which means that the function is concave and the maximum is unique. Therefore, $p=q$ is the only maximizer and the Brier scoring rule for multiclass classification is strictly proper.
A.8 Proof that the Focal Calibration Loss is Strictly Proper
Proof.
We aim to show that $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ is minimized if and only if $\hat{\boldsymbol{p}}=\boldsymbol{y}$ when $\lambda>0$ and $\gamma>0$, and that it is not strictly proper when $\lambda=0$.
Case 1: $\lambda>0$
Let $\boldsymbol{y}\in\Delta^{K}$ be the true probability vector, and $\hat{\boldsymbol{p}}\in\Delta^{K}$ be the predicted probability vector. We consider two scenarios:
(a) When $\hat{\boldsymbol{p}}=\boldsymbol{y}$.
At $\hat{\boldsymbol{p}}=\boldsymbol{y}$, the calibration loss is:
$\mathcal{L}_{\text{calib}}(\boldsymbol{y},\boldsymbol{y})=\left\|\boldsymbol{y}-\boldsymbol{y}\right\|_{2}^{2}=0.$ (74)
Assuming $\boldsymbol{y}$ is a one-hot vector with $y_{j}=1$ for the true class $j$, the focal loss becomes:
$\mathcal{L}_{\text{focal}}^{\gamma}(\boldsymbol{y},\boldsymbol{y})=-(1-y_{j})^{\gamma}\log y_{j}=-0^{\gamma}\cdot 0=0.$ (75)
Thus, the total loss is:
$\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{y},\boldsymbol{y})=0+\lambda\cdot 0=0.$ (76)
(b) When $\hat{\boldsymbol{p}}\neq\boldsymbol{y}$.
The calibration loss is strictly positive:
$\mathcal{L}_{\text{calib}}(\hat{\boldsymbol{p}},\boldsymbol{y})=\left\|\hat{\boldsymbol{p}}-\boldsymbol{y}\right\|_{2}^{2}>0.$ (77)
The focal loss for the true class $j$ satisfies:
$\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\boldsymbol{p}},\boldsymbol{y})=-(1-\hat{p}_{j})^{\gamma}\log\hat{p}_{j}>0\quad\text{for }\hat{p}_{j}<1.$ (78)
Thus, the total loss is:
$\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\hat{\boldsymbol{p}},\boldsymbol{y})=\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\boldsymbol{p}},\boldsymbol{y})+\lambda\,\mathcal{L}_{\text{calib}}(\hat{\boldsymbol{p}},\boldsymbol{y})>0.$ (79)
Conclusion for $\lambda>0$:
Since the loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ attains its unique global minimum of zero when $\hat{\boldsymbol{p}}=\boldsymbol{y}$, it is strictly proper when $\lambda>0$.
Case 2: $\lambda=0$
When $\lambda=0$, the FCL reduces to the focal loss:
$\mathcal{L}_{\text{FCL}}^{\gamma,0}(\hat{\boldsymbol{p}},\boldsymbol{y})=\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\boldsymbol{p}},\boldsymbol{y}).$ (80)
The focal loss is not strictly proper because:
(a) Multiple Minimizers:
Consider any $\hat{\boldsymbol{p}}$ where $\hat{p}_{j}=1$ and $\hat{p}_{k}=0$ for $k\neq j$. In this case:
$\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\boldsymbol{p}},\boldsymbol{y})=0.$ (81)
However, suppose $\hat{p}_{j}=q<1$ but still significantly higher than other $\hat{p}_{k}$. Depending on $\gamma$, the focal loss may not penalize this deviation strongly, and $\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\boldsymbol{p}},\boldsymbol{y})$ may still be close to zero. Therefore, the minimum loss may not be uniquely attained at $\hat{\boldsymbol{p}}=\boldsymbol{y}$.
(b) Lack of Strict Convexity:
The focal loss is not strictly convex with respect to $\hat{\boldsymbol{p}}$, especially when $\gamma=0$, where it reduces to the cross-entropy loss, which is proper but not strictly proper without additional regularization.
Conclusion for $\lambda=0$:
Since the minimum of $\mathcal{L}_{\text{FCL}}^{\gamma,0}$ is not uniquely achieved at $\hat{\boldsymbol{p}}=\boldsymbol{y}$, the loss is not strictly proper when $\lambda=0$.
Overall Conclusion
The Focal Calibration Loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ is strictly proper if and only if $\lambda>0$ and $\gamma>0$. ∎
A.9 Focal Loss Relative with Entropy Regularised KL Divergence
This section we illustrate why focal loss tends to favor accurate yet less confident solutions. We demonstrate that it inherently balances minimizing the KL-divergence and maximizing the entropy based on the parameter $\gamma$. Let $\mathcal{L}_{f}$ represent the focal loss with parameter $\gamma$, and $\mathcal{L}_{c}$ represent the cross-entropy loss between $\hat{p}$ and $q$. Here, $K$ denotes the number of classes, and $q_{k}$ and $\hat{p}_{k}$ represent the ground-truth and predicted probabilities for the $k$-th class, respectively. We consider the following straightforward extension of focal loss:
$\mathcal{L}_{\text{focal}}(\hat{\boldsymbol{p}},\boldsymbol{y})$ $=-\sum_{k=1}^{K}y_{ik}(1-\hat{p}_{ik})^{\gamma}\log\hat{p}_{ik}$ (82)
$\geq-\sum_{k=1}^{K}y_{ik}(1-\gamma\hat{p}_{ik})\log\hat{p}_{ik}$
$\text{By Bernoulli's inequality}\,\forall\gamma\geq 1,\text{since}\,\hat{p}_{ik}\in[0,1]$
$=-\sum_{k=1}^{K}y_{ik}\log\hat{p}_{ik}-\gamma\left|\sum_{k=1}^{K}y_{ik}\hat{p}_{ik}\log\hat{p}_{ik}\right|\quad\forall k,\log\hat{p}_{ik}\leq 0$
$\geq-\sum_{k=1}^{K}y_{ik}\log\hat{p}_{ik}+\gamma\sum_{k=1}^{K}\hat{p}_{ik}\log\hat{p}_{ik}=\mathcal{L}_{c}-\gamma H[\hat{p}$]
[…]
Given that $\mathcal{L}_{c}=KL(q\|\hat{p})+H[q]$, combining it to inequality Eq. (82):
$\mathcal{L}_{\text{focal}}(\hat{\boldsymbol{p}},\boldsymbol{y})\geq KL(q\|\hat{\boldsymbol{p}})+H[q]-\gamma H[\hat{\boldsymbol{p}}$] (83)
In the case of one-hot encoding (Delta distribution for $q$), focal loss will maximize $\hat{p}_{k}\log\hat{p}_{k}$ (let $k$ be the ground-truth class index), the component of the entropy of $\hat{\boldsymbol{p}}$ corresponding to the ground-truth index. Thus, it will prefer learning $\hat{p}$ such that $\hat{p}_{k}$ is assigned a higher value (because of the KL term), but not too high (because of the entropy term), and will ultimately avoid preferring overconfident models (by contrast to cross-entropy loss).
Figure 6: Optimal Predicted Probability ($\hat{p_{ik}}$) for Various Values of Ground True Probability ($y_{ik}$) with Calibration Analysis. Circular markers with ”+” and ”-” signs represent the areas where the calibration term increases or decreases in visualizing how it adjust the predicted probabilities.
A.10 Hyperparameter Selection for Focal Calibration Loss
Table 5: $\gamma$ and $\lambda$ Selection
Dataset Model $\gamma$ $\lambda$
CIFAR-100 ResNet-50 2.0 0.5
ResNet-110 4.0 1.5
Wide-ResNet-26-10 3.0 0.5
DenseNet-121 3.0 0.5
CIFAR-10 ResNet-50 4.0 1.5
ResNet-110 4.0 1.5
Wide-ResNet-26-10 2.0 0.5
DenseNet-121 2.0 0.5
Tiny-ImageNet ResNet-50 3.0 0.5
20 Newsgroups Global Pooling CNN 5.0 4.5
AG News Global Pooling CNN 4.0 3.0
FinSen Global Pooling CNN 4.0 4.0
A.11 Decision Boundary on Focal Loss and Focal Calibration Loss (FCL)
We compared the decision boundaries of a neural network trained with Focal Loss (FL) and Focal Calibration Loss (FCL) on a noisy, hard-to-classify synthetic dataset, using the same $\gamma$ = 10 for both models. The dataset, generated with noise=0.2, created overlapping classes that are difficult to separate. Both models were trained using a simple feedforward neural network with two hidden layers of 10 units and optimized with Adam for 500 epochs. While the FL model generated sharp decision boundaries focusing on hard samples, the FCL model, with $\lambda$ = 1.5, produced a smoother boundary as depicted in Fig. 7, effectively balancing accuracy and calibration.
Figure 7: Decision boundary comparison between Focal Loss (left) and Focal Calibration Loss (FCL) (right) on a synthetic dataset with hard samples. In both models, blue and red regions represent the areas classified as Class 0 and Class 1, respectively. The yellow boxes highlight regions where hard samples from both classes overlap, and the green boxes indicate areas where FCL provides a smoother decision boundary compared to Focal Loss.
A.12 Table 6 AdaECE Evaluation Result
Table 6: $\downarrow$ AdaECE (%) for various methods both before (pre) and after (post) applying temperature scaling.
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T
CIFAR-100 ResNet-50 17.99 3.38(2.2) 5.46 4.24(1.1) 15.05 3.42(1.9) 6.72 5.37(1.1) 5.64 2.84(1.1) 8.79 2.27(1.3) 3.74 2.11(1.1)
ResNet-110 19.28 6.27(2.3) 6.51 3.75(1.2) 18.83 4.86(2.3) 9.68 8.11(1.3) 10.90 4.13(1.3) 11.64 4.48(1.3) 4.88 3.27(1.1)
Wide-ResNet-26-10 15.16 3.23(2.1) 4.08 3.11(1.1) 13.55 3.83(2.0) 3.73 3.73(1.0) 2.40 2.38(1.1) 5.36 2.38(1.2) 2.35 2.35(1.0)
DenseNet-121 19.07 3.82(2.2) 3.92 2.41(1.1) 17.37 3.07(2.0) 8.62 5.92(1.1) 3.35 1.80(1.1) 6.69 1.69(1.2) 1.47 1.47(1.0)
CIFAR-10 ResNet-50 4.22 2.11(2.5) 1.85 1.34(1.1) 4.67 2.01(2.6) 4.28 3.20(0.9) 1.64 1.64(1.0) 1.28 1.28(1.0) 0.75 0.75(1.0)
ResNet-110 4.78 2.42(2.6) 2.52 1.72(1.2) 5.21 2.66(2.8) 4.57 3.62(0.9) 1.76 1.32(1.1) 1.69 1.42(1.1) 1.25 1.25(1.0)
Wide-ResNet-26-10 3.22 1.62(2.2) 1.94 1.94(1.0) 3.58 1.83(2.2) 4.58 2.55(0.8) 1.84 1.63(0.9) 3.16 1.20(0.8) 1.39 1.39(1.0)
DenseNet-121 4.69 2.28(2.4) 1.84 1.84(1.0) 4.97 2.69(2.4) 4.60 3.36(0.9) 1.58 1.62(0.9) 0.79 1.32(0.9) 1.28 1.28(1.0)
Tiny-ImageNet ResNet-50 16.02 4.99(1.5) 5.00 3.60(0.9) 13.15 4.73(1.3) 15.05 5.36(0.7) 1.31 1.31(1.0) 2.66 2.66(1.0) 10.22 1.34(0.8)
NLP 20 Newsgroups Global Pooling CNN 17.91 2.23(3.4) 13.57 3.11(2.3) 15.21 6.47(2.2) 4.39 2.63(1.1) 8.65 3.78(1.5) 18.44 3.81(2.3) 15.29 1.48(2.2)
NLP AG News Global Pooling CNN 4.77 0.67(2.0) 6.78 2.77(2.8) 1.80 1.78(1.1) 3.84 1.70(0.8) 6.69 0.57(3.0) 2.99 2.51(0.8) 0.97 0.38(1.1)
NLP FinSen Global Pooling CNN 0.91 0.24(2.5) 0.35 0.35(1.0) 1.23 0.37(0.5) 4.39 4.5(0.6) 0.28 0.43(0.9) 0.23 0.43(0.7) 0.46 0.24(1.3)
A.13 Table 7 Classwise-ECE Evaluation Result
Table 7: $\downarrow$ Classwise-ECE (%) for various methods both before (pre) and after (post) applying temperature scaling.
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T
CIFAR-100 ResNet-50 0.39 0.22(2.2) 0.21 0.22(1.1) 0.34 0.22(1.9) 0.21 0.22(1.1) 0.21 0.20(1.1) 0.24 0.22(1.3) 0.19 0.19(1.1)
ResNet-110 0.42 0.21(2.3) 0.22 0.23(1.2) 0.41 0.22(2.3) 0.25 0.23(1.3) 0.27 0.22(1.3) 0.28 0.21(1.3) 0.20 0.21(1.1)
Wide-ResNet-26-10 0.34 0.21(2.1) 0.19 0.20(1.1) 0.31 0.21(2.0) 0.20 0.20(1.0) 0.18 0.20(1.1) 0.19 0.20(1.2) 0.20 0.20(1.0)
DenseNet-121 0.42 0.22(2.2) 0.21 0.21(1.1) 0.39 0.23(2.0) 0.23 0.21(1.1) 0.20 0.21(1.1) 0.22 0.21(1.2) 0.19 0.19(1.0)
CIFAR-10 ResNet-50 0.87 0.37(2.5) 0.46 0.39(1.1) 0.97 0.55(2.6) 0.80 0.54(0.9) 0.41 0.41(1.0) 0.45 0.45(1.0) 0.36 0.36(1.0)
ResNet-110 1.00 0.54(2.6) 0.55 0.46(1.2) 1.08 0.60(2.8) 0.75 0.50(0.9) 0.48 0.46(1.1) 0.46 0.52(1.1) 0.38 0.38(1.0)
Wide-ResNet-26-10 0.68 0.34(2.2) 0.37 0.37(1.0) 0.77 0.41(2.2) 0.95 0.37(0.8) 0.44 0.34(0.9) 0.82 0.33(0.8) 0.33 0.33(1.0)
DenseNet-121 0.98 0.54(2.4) 0.43 0.43(1.0) 1.02 0.53(2.4) 0.75 0.48(0.9) 0.43 0.41(0.9) 0.40 0.41(0.9) 0.37 0.37(1.0)
Tiny-ImageNet ResNet-50 0.23 0.17(1.5) 0.17 0.16(0.9) 0.21 0.17(1.3) 0.21 0.17(0.7) 0.16 0.16(1.0) 0.16 0.16(1.0) 0.19 0.16(0.8)
NLP 20 Newsgroups Global Pooling CNN 1.95 0.83(3.4) 1.56 0.82(2.3) 1.77 1.10(2.2) 0.93 0.91(1.1) 1.40 1.19(1.5) 2.01 1.04(2.3) 1.71 0.80(2.2)
NLP AG News Global Pooling CNN 2.49 0.86(2.0) 3.46 1.64(2.8) 1.07 0.96(1.1) 2.69 1.57(0.8) 3.39 0.77(3.0) 2.43 1.92(0.8) 0.86 1.02(1.1)
NLP FinSen Global Pooling CNN 0.18 0.16(2.5) 0.17 0.17(1.0) 0.32 0.19(0.5) 0.82 0.18(0.6) 0.15 0.16(0.9) 0.17 0.16(0.7) 0.15 0.15(1.3)
A.14 Proof of Thm. (4.5): Post-Processing Gap of focal calibration loss and focal loss
Proof.
Let the post-processed prediction be defined as $\kappa(\hat{p}_{ik})=\hat{p}_{ik}+\eta(\hat{p}_{ik})$, where $\eta(\hat{p}_{ik})$ is the post-processing adjustment function applied to the predicted probability $\hat{p}_{ik}$. This gives the post-processed version of the focal calibration loss (FCL) as:
$\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\kappa(\hat{\boldsymbol{p}}),\boldsymbol{y})=\mathcal{L}_{\text{focal}}(\kappa(\hat{\boldsymbol{p}}),\boldsymbol{y})+\lambda\mathcal{L}_{\text{calib}}(\kappa(\hat{\boldsymbol{p}}),\boldsymbol{y})$ (84)
where $\mathcal{L}_{\text{focal}}$ is the focal loss, $\mathcal{L}_{\text{calib}}$ is the calibration loss, and $\lambda\geq 0$ is a weighting parameter controlling the calibration term.
Following Definition 2.3, the post-processing gap for the FCL $\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}}^{\gamma,\lambda})$is defined as:
$\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(f(x),y)]-\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\kappa(f(x)),y)]$ (85)
Step 1: Expanding the Post-Processing Gap for FCL
Substituting the definition of the focal calibration loss, we obtain:
$\begin{split}\text{pGap}_{\mathcal{D}}\left(\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}\right)=\ &\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{focal}}\left(f(x),y\right)\right]\\ &+\lambda\,\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{calib}}\left(f(x),y\right)\right]\\ &-\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{focal}}\left(\kappa\left(f(x)\right),y\right)\right]\\ &-\lambda\,\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{calib}}\left(\kappa\left(f(x)\right),y\right)\right]\end{split}$ (86)
Step 2: Special Case when $\lambda=0$
When $\lambda=0$, the focal calibration loss simplifies to the focal loss, as the calibration term is removed. Thus, the post-processing gap for FCL reduces to the post-processing gap for FL:
$\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}}^{\gamma,\lambda=0})=\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{focal}})$ (87)
Step 3: General Case when $\lambda>0$
When $\lambda>0$, the calibration term influences the post-processing gap by helping to reduce the calibration error. We focus on the term:
$\lambda\left(\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\text{calib}}(f(x),y)]-\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\text{calib}}(\kappa(f(x)),y)]\right)$ (88)
This term measures the improvement in calibration achieved through post-processing $\kappa$. The calibration loss, $\mathcal{L}_{\text{calib}}(f(x),y)$, can be defined as a squared loss: $\mathcal{L}_{\text{calib}}(\hat{\boldsymbol{p}},\boldsymbol{y})=\sum_{k=1}^{K}(y_{ik}-\hat{p}_{ik})^{2}.$
Thus, the calibration gap $\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{calib}})$ becomes:
$\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}(y_{ik}-\hat{p}_{ik})^{2}\right]-\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}(y_{ik}-\kappa(\hat{p}_{ik}))^{2}\right]$ (89)
$=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}(y_{ik}-\hat{p}_{ik})^{2}-\left(y_{ik}-\hat{p}_{ik}-\eta(\hat{p}_{ik})\right)^{2}\right]$
Step 4: Simplifying $\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{calib}})$
Expanding the square term, we get:
$\begin{aligned} &\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}\left((y_{ik}-\hat{p}_{ik})^{2}-\left((y_{ik}-\hat{p}_{ik})^{2}-2(y_{ik}-\hat{p}_{ik})\eta(\hat{p}_{ik})+\eta(\hat{p}_{ik})^{2}\right)\right)\right]\\ &=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}2(y_{ik}-\hat{p}_{ik})\eta(\hat{p}_{ik})-\eta(\hat{p}_{ik})^{2}\right]\end{aligned}$
Thus, the total post-processing gap becomes:
$\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}}^{\gamma,\lambda})$ $=\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{focal}})+\lambda\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{calib}})$ (91)
$\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{calib}})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}2\underbrace{(y_{ik}-\hat{p}_{ik})\eta(\hat{p}_{ik})}_{\textbf{CalibError}}-\eta(\hat{p}_{ik})^{2}\right]$
Step 5: Analyzing Overconfidence (OC) Cases
When $\hat{p}_{ik}>y_{ik}$, $\eta(\hat{p}_{ik})>0$. In this case, the term $2(y_{ik}-\hat{p}_{ik})\eta(\hat{p}_{ik})$ is negative, and $\eta(\hat{p}_{ik})^{2}<0$, which means that the calibration term overall is negative. This reduction in the post-processing gap brings the prediction closer to the true label, reducing the pGap.
Step 6: Analyzing Underconfidence (UC) Case
When $\hat{p}_{ik}<y_{ik}$, $\eta(\hat{p}_{ik})>0$, and thus $2(y_{ik}-\hat{p}_{ik})\eta(\hat{p}_{ik})$ is positive. This positive term increases the predicted value, reducing the gap between $\hat{p}_{ik}$ and $y_{ik}$. Additionally, the term $\eta(\hat{p}_{ik})^{2}$ is still negative but smaller in magnitude, ensuring that the gap is reduced.
Since $\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}2(y_{ik}-\hat{p}_{ik})\eta(\hat{p}_{ik})-\eta(\hat{p}_{ik})^{2}\right]$ ensures that pGap is not increased, maintaining or reducing the gap, as depicted in Fig. 6, we conclude that:
$\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}})\leq\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{focal}}).$ (92)
A.15 Ablation Study: L1 vs. L2 Loss
We did not present results for L1 or other $L_{p}$ losses in the main text because our preliminary experiments indicated that they perform worse than L2 in terms of calibration. Table 8 shows example comparisons (test error and ECE) using Mean Absolute Error (L1) vs. Mean Squared Error (L2) on CIFAR-10 and CIFAR-100. We observe that L2 consistently achieves lower ECE, so we adopt L2 (MSE) for our focal calibration loss. As shown, L2 (MSE) outperforms L1 (MAE) in terms of ECE across different architectures, which motivated our choice of L2 in all experiments.
Table 8: Comparing L1 (MAE) vs. L2 (MSE) for calibration on CIFAR-10 and CIFAR-100.
Dataset Model Loss Test Error (%) ECE (%)
CIFAR-100 ResNet-110 L1 22.38 4.45
ResNet-110 L2 21.02 4.19
WideResNet L1 19.99 4.82
WideResNet L2 19.81 2.22
DenseNet-121 L1 21.66 3.42
DenseNet-121 L2 21.80 1.31
ResNet-50 L1 21.64 5.65
ResNet-50 L2 21.81 3.71
CIFAR-10 ResNet-110 L1 5.09 1.46
ResNet-110 L2 4.75 0.95
WideResNet L1 4.12 2.44
WideResNet L2 4.16 0.92
DenseNet-121 L1 5.37 4.00
DenseNet-121 L2 4.59 0.66
ResNet-50 L1 5.08 1.78
ResNet-50 L2 4.56 0.76
A.16 CheXNet: Interpretability and Per-Class AUROC Results
This section provides an extended discussion on how FCL improves interpretability and performance in medical imaging tasks, specifically for Chest X-ray anomaly detection (CheXNet).
Interpretability with Grad-CAM Heatmaps. Grad-CAM heatmaps offer a widely accepted visualization tool for highlighting the most influential regions of a model’s activation. By adopting FCL in training, medical practitioners gain a more intuitive understanding of the model’s decision-making process, thereby enhancing trust and transparency. As illustrated in Fig. 1 (in the main text), FCL produces clearer and more precise anomaly-localization maps compared to baseline methods. This improvement not only has practical utility in diagnostic tasks but also provides stronger interpretability, crucial for the stringent reliability standards in medical applications. Consistent with Section 5.3.1, FCL achieves superior calibration metrics while delivering finer visual explanations—bridging the gap between purely numerical evaluations and real-world clinical needs.
Training Procedure. We followed the training procedure outlined by Rajpurkar et al. (Rajpurkar, 2017), including the ten-crop augmentation for both validation and test phases. Table 9 compares per-class AUROC scores for CheXNet using BCE vs. our FCL (along with earlier baselines). Note that FCL consistently yields strong performance gains, particularly in terms of average AUROC and reliability scores (ECE, MCE).
Table 9: AUROC comparison (%) over previous SOTA results on the ChestX-ray14 dataset. ECE* and MCE† are shown in parentheses.
Pathology Wang et al.  (Wang et al., 2017) Yao et al.  (Yao et al., 2017) CheXNet (BCE)  (Rajpurkar, 2017) FCL (Ours)
Atelectasis 0.716 0.772 0.8094 0.8332
Cardiomegaly 0.807 0.904 0.9248 0.9156
Effusion 0.784 0.859 0.8638 0.8862
Infiltration 0.609 0.695 0.7345 0.7151
Mass 0.706 0.792 0.8676 0.8701
Nodule 0.671 0.717 0.7802 0.8017
Pneumonia 0.633 0.713 0.7680 0.7813
Pneumothorax 0.806 0.841 0.8887 0.8866
Consolidation 0.708 0.788 0.7901 0.8192
Edema 0.835 0.882 0.8878 0.8996
Emphysema 0.815 0.829 0.9371 0.9417
Fibrosis 0.769 0.767 0.8047 0.8559
Pleural Thickening 0.708 0.765 0.8062 0.7958
Hernia 0.767 0.914 0.9164 0.9370
Average 0.742 (/) 0.798 (/) 0.8440 (0.1738*, 0.6134†) 0.8527 (0.1292*, 0.3248†)
Overall, FCL consistently enhances both calibration quality and interpretability in CheXNet. These improvements align with the high standards of reliability and transparency demanded in clinical practice, making FCL particularly suitable for real-world medical diagnostic workflows.
Figure 8: Reliability Diagram of Various Methods before Temperature Scaling (CIFAR10, ResNet-50, 100 - 300 Epochs). The top/middle/bottom row corresponds to the results at 100/200/300 epochs. Each row shows the performance of Cross Entropy, FLSD-53, Dual Focal Loss, and Focal Calibration Loss.
A.17 Reliability Diagram and ROC Diagram
Figure 9: Reliability Diagram of Various Methods before Temperature Scaling(CIFAR10, ResNet-110, 300 Epochs).
(a) ResNet-50 (pre-T)
(b) ResNet-110 (pre-T)
(c) Wide-ResNet (pre-T)
(d) DenseNet (pre-T)
(e) ResNet-50 (post-T)
(f) ResNet-110 (post-T)
(g) Wide-ResNet (post-T)
(h) DenseNet (post-T)
Figure 10: ROC plots obtained from different architectures trained on CIFAR-10 and tested on SVHN.
(a) ResNet-50 (pre-T)
(b) ResNet-110 (pre-T)
(c) Wide-ResNet (pre-T)
(d) DenseNet (pre-T)
(e) ResNet-50 (post-T)
(f) ResNet-110 (post-T)
(g) Wide-ResNet (post-T)
(h) DenseNet (post-T)
Figure 11: ROC plots obtained from different architectures trained on CIFAR-10 and tested on CIFAR-10-C (OoD).
Table 10: $\downarrow$ ECE (%) for various methods both before (pre) and after (post) applying temperature scaling (Bin = 15) 44footnotemark: 4.
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T
CIFAR-100 $\text{ViT}_{224}$ 13.11 3.30(1.3) 3.66 3.23(1.1) 11.52 3.34(1.3) 3.83 3.24(1.1) 5.33 3.46(1.1) 3.42 3.42(1.0) 3.72 2.57(1.1)
$\text{DeiT}_{\text{small}}$ 7.46 3.54(1.1) 2.17 2.17(1.0) 7.88 2.55(1.2) 2.75 2.75(1.0) 3.27 3.27(1.0) 1.93 1.93(1.0) 1.75 1.75(1.0)
CIFAR-10 $\text{ViT}_{224}$ 7.39 1.68(1.3) 1.87 1.87(1.0) 3.47 1.46(1.1) 1.79 1.79(1.0) 8.70 2.11(0.7) 3.71 2.15(0.9) 1.36 1.36(1.0)
$\text{DeiT}_{\text{small}}$ 4.42 2.52(1.2) 2.02 2.02 (1.0) 6.01 1.78(1.2) 1.84 1.84(1.0) 10.65 2.79(0.7) 5.11 2.79(0.8) 1.47 1.47(1.0)